### v1

# Promptus: Can Prompts Streaming Replace Video Streaming with Stable Diffusion

Abstract
With the exponential growth of video traffic, traditional video streaming systems are approaching their limits in compression efficiency and communication capacity. To further reduce bitrate while maintaining quality, we propose Promptus, a disruptive novel system that streaming prompts instead of video content with Stable Diffusion, which converts video frames into a series of "prompts" for delivery. To ensure pixel alignment, a gradient descent-based prompt fitting framework is proposed. To achieve adaptive bitrate for prompts, a low-rank decomposition-based bitrate control algorithm is introduced. For inter-frame compression of prompts, a temporal smoothing-based prompt interpolation algorithm is proposed. Evaluations across various video domains and real network traces demonstrate Promptus can enhance the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Moreover, Promptus achieves real-time video generation from prompts at over 150 FPS. To the best of our knowledge, Promptus is the first attempt to replace video codecs with prompt inversion and the first to use prompt streaming instead of video streaming. Our work opens up a new paradigm for efficient video communication beyond the Shannon limit.
1 Introduction
With the rapid development of streaming applications (such as video-on-demand [2, 8], live video [14, 7], video conferencing [15, 11], cloud gaming [3, 13], etc.), the traffic of network video has been continuously growing. To reduce traffic, video codecs represented by VP8 [16], VP9 [33], H.264 [4] and H.265 [5] are widely used to compress videos. These codecs achieve compression by removing spatial and temporal redundancies. However, these redundancies are limited, so there is an upper bound on the compression ratio (subject to the Shannon limit [40]). In order to further compress the video, non-redundant content in the video will be discarded, which will greatly degrade the video quality, such as causing blurring and blocking artifacts.footnotetext: Promptus will be released at https://pku-netvideo.github.io/ once published.
Figure 1: Promptus can invert a given image into a prompt. Based on this prompt, Stable Diffusion can generate an almost identical image to the original. In contrast, existing methods can only generate semantically similar images.
To improve compression efficiency, some deep learning-based codecs and streaming frameworks have been proposed in recent years. Neural-embedded codecs [31, 30, 19, 37, 47] replace handcrafted modules (such as motion vector estimation [45]) in traditional codec frameworks with neural networks (such as autoencoders), learning more intelligent and efficient compression capabilities through end-to-end optimization. However, due to still following the traditional encoding framework, the improvement in compression ratio is limited. Neural-enhanced streaming [50, 49, 26, 52, 56] transmits distorted low-bitrate videos and then restores high-fidelity details using post-processing algorithms such as super-resolution. However, these post-processing algorithms rely on prior knowledge learned from training sets. Due to the domain gap, their performance will degrade in unseen scenes [48]. Generative streaming [24, 44, 42, 29] transmits small-sized semantic information, and the receiver can generate video based on these semantics. However, most generative algorithms are designed for specific tasks (e.g., talking video generation driven by facial keypoints [24, 44]), limiting their generality.
With the popularity of AIGC (AI-generated content), Stable Diffusion [38, 10, 9] has attracted extensive attention thanks to its powerful text-to-image generation capability. By pre-training on an internet-scale dataset [6], Stable Diffusion learns prior knowledge of all human visual domains, and simultaneously learns the mapping from text to images. Therefore, Stable Diffusion can generate high-fidelity images based on a brief prompt composed of a few words.
We are motivated by this question: is it possible for Stable Diffusion to replace the video codecs? During streaming, the sender streams prompts instead of streaming encoded videos and the receiver generates videos instead of decoding videos. In this way, the traffic of network video is reduced from the video scale to the text scale, achieving communication efficiency beyond the Shannon limit.
In this paper, we propose Promptus, a system that inverts video frames into Stable Diffusion prompts, achieving ultra-low bitrate video streaming. To bring this vision to fruition, we address the following technical challenges:
First, how to ensure pixel alignment between the generated frames and the real frames. To invert a frame into a prompt, the most straightforward and powerful approach is to manually write a textual description. However, the frame generated using this description can only guarantee semantic consistency with the original frame, while the differences at the pixel level are often substantial, as shown in Figure 1. To achieve pixel alignment, Promptus proposes a gradient descent-based prompt fitting framework. Specifically, the prompt is randomly initialized and then used to generate frame. The pixel-wise loss between the generated frame and the real frame is calculated. The partial derivative of the loss with respect to the prompt will be calculated and the prompt is iteratively optimized using gradient descent. To implement this framework, Promptus employs single-step denoising instead of iterative denoising to avoid higher-order derivatives. Second, Promptus uses embeddings as prompts instead of text to avoid non-differentiability. Third, Promptus uses a noisy previous frame instead of random noise to reduce latent space distance. Fourth, Promptus combines reconstruction and perceptual loss to enhance the perceptual quality of the generated frames.
Second, how to control the bitrate of the prompt. Since Promptus uses embeddings as prompts, which are matrices with fixed dimensions, the bitrate of the prompt cannot adapt to dynamic network bandwidth. To address this, Promptus proposes a low-rank bitrate control algorithm. Specifically, Promptus integrates the inverse process of low-rank decomposition into the gradient descent, directly fitting the decomposed prompt. The rank is used to control the trade-off between the quality and bitrate of the prompt. When the rank is higher, the representational capability of the prompt is better, and it can describe more details in the frame, but the data size is also larger. Therefore, Promptus adaptively selects the prompt rank based on the currently available bandwidth.
Third, how to perform inter-frame compression on prompts. Promptus inverts each frame into an independent prompt without considering the correlation between prompts across frames. Therefore, when streaming video, Promptus needs to transmit prompts for each frame, resulting in the bitrate increasing linearly with the frame rate. To address this, Promptus adds a temporal smoothing regularization during prompt fitting, ensuring that temporally close frames are also sufficiently close in the prompt space. With this, Promptus only needs to sparsely transmit prompts for a few keyframes, while the prompts for the remaining frames can be approximated through linear interpolation of the keyframe prompts.
We evaluated Promptus on test videos from different domains and under real network traces. The results show that: First, Promptus achieves scalable bitrate, and its quality advantage over the baselines becomes more significant as the target bitrate decreases. Second, Promptus is general to different video domains, and the more complex the video content, the greater the quality advantage of Promptus compared to the baselines. Third, under real-world network traces, Promptus enhances the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE [27, 38] and H.265 [5], respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Fourth, Promptus can generate videos from prompts in real-time at a speed exceeding 150 FPS.
The contributions of this paper are summarized as follows: (1) We propose Promptus, which, to the best of our knowledge, is the first attempt to replace video codecs with prompt inversion and also the first to use prompt streaming to replace video streaming (§3). (2) We propose a gradient descent-based prompt fitting framework, achieving pixel-aligned prompt inversion for the first time (§3.1). (3) We build a video streaming system based on Promptus (§4).
2 Motivation and related work
2.1 Video codec and streaming
Traditional codecs. In network video traffic, most of the content is encoded using traditional codecs represented by VP8 [16], VP9 [33], H.264 [4] and H.265 [5]. These traditional video codecs achieve compression primarily by eliminating redundancy in the video. Specifically, the codecs exploit the correlation between adjacent pixel blocks to reduce spatial redundancy in intra-frame prediction. Besides, considering the similarity between consecutive video frames, the codecs employ inter-frame prediction techniques to eliminate temporal redundancy. For moving objects, the codecs perform motion estimation and compensation to further reduce redundancy. Although these techniques achieve efficient compression, the compression ratio has an upper limit due to the finite amount of redundancy. Further compression inevitably requires discarding some non-redundant information, resulting in a drastic decrease in video quality. For example, as shown in Figure 11, when the bitrate is extremely low, videos compressed using H.265 exhibit significant blurriness.
Neural-embedded codecs. With the development of deep learning, handcrafted modules in traditional codecs are being replaced by neural networks [31, 30, 19, 37, 47]. Compared to handcrafted modules, these embedded neural networks can learn more complex and nonlinear motion patterns and intra-frame correlations, thereby achieving lower distortion at the same bitrate. Furthermore, due to the ability to optimize end-to-end, neural-embedded codecs can be trained for tasks beyond compression, such as loss resilience. However, since neural-embedded codecs still adhere to the traditional coding framework, aiming to fully remove intra-frame or inter-frame redundancy, the improvement in compression ratio is limited.
Neural-enhanced streaming. In contrast to reducing redundancy, neural-enhanced streaming [50, 49, 26, 52, 56] actively discards most of the information (including non-redundant information) and transmits highly distorted low-bitrate videos. Then, at the receiver side, neural network-based post-processing algorithms (such as super-resolution) are used to restore high-fidelity details. Neural-enhanced streaming leverages the powerful image restoration and enhancement capabilities of neural networks, effectively recovering lost textures and details, thus ensuring video quality while achieving substantial compression. However, these post-processing algorithms rely on prior knowledge learned from training datasets, such as texture features and edge structures. Due to the complex and diverse nature of real-world video, there exists domain gaps with the training set. The performance of these algorithms often degrades in unseen scenarios [48].
Generative streaming. Compared to using neural networks for post-processing, generative streaming directly uses neural networks to generate videos to handle higher compression ratios [24, 44, 42, 29]. For example, in the context of video conferencing, Face-vid2vid [44, 24] extracts facial keypoints in real-time at the sender side for transmission. At the receiver side, it generates dynamic facial videos using the keypoints and static facial images. The bitrate of facial keypoints is much lower than that of videos, thus compressing video conferences to extremely low bitrates. Similarly, Gemino [42] transmits video streams at extremely low resolution. At the receiver side, it estimates facial motion fields from the low-resolution video, then utilizes the motion fields and high-resolution facial images to generate high-resolution facial videos. Instead of driving static images, Reparo [29] proposes a token-based generative streaming. It first uses VQGAN [20] to train a codebook of facial visual features, then maps the video to latent variables using VAE [27], and finally quantizes the latent variables into tokens according to the codebook. Since tokens are indices of the codebook, the bitrate is extremely low. In general, generative streaming can greatly compress videos, but it is often designed for specific tasks (such as video conferencing) and lacks generality.
2.2 Stable diffusion
Stable Diffusion [38, 10, 9] is the most popular open-source text-to-image generative model. It learns a denoising process from 5.85 billion image-text pairs [6], enabling it to generate high-quality images by denoising the pure noise images. Different noise images lead to different generated images. Since the noise images are randomly sampled from Gaussian distribution, the images generated by Stable Diffusion are random and uncontrollable. Therefore, to control the content of the generated images, Stable Diffusion also receives user-input prompts (in natural language) as the condition for denoising, thereby generating images that align with the semantic descriptions of the prompts. Specifically, as shown in Figure 2, $p$ represents the user-input prompt. Stable Diffusion first performs word embedding and semantic extraction using the CLIP model [36], converting discrete natural language into a continuous text embedding $c$ (an $m*n$ matrix). The text embeddings and the randomly sampled noise image $N$ are input into the denoising process of Stable Diffusion, generating the denoised $N^{\prime}$. Next, $N^{\prime}$ is input into the denoising process again for $T$ iterations to obtain the denoised $Z$. Since the denoising process occurs in the latent space, $Z$ needs to be input into the VAE Decoder [27] to generate the image $x$ in the pixel space.
Figure 2: How does Stable Diffusion generate high-quality images from text prompts. The details are elaborated in §2.2.
Despite the impressive image generation capabilities of Stable Diffusion, its potential in the field of video streaming has been overlooked. However, we believe that it can significantly enhance the performance of video streaming for the following reasons:
  • Stable Diffusion can effectively reduce the bandwidth overhead of videos. It can generate high-quality images using brief prompts. As shown in Figure 2, a prompt consisting of only a few words can enable Stable Diffusion to generate images with high resolution and complex textures. In codec-based streaming, such high-quality images require high bitrates. Instead of transmitting image, if we solely transmit prompts and generate images at the receiving end, the bitrate can be reduced from the image scale to the prompt scale, greatly improving communication efficiency.
  • Stable Diffusion is generalizable to most video domains. Trained on an internet-scale dataset [6] (5.85 billion images), Stable Diffusion has learned priors for nearly all video domains present on the internet. Consequently, it does not suffer from the domain gap issue and can generate images of any style and content. This gives Stable Diffusion the potential to become a versatile paradigm for video streaming.
  • Stable Diffusion can generate images in real-time. With ongoing advancements in Stable Diffusion, variants capable of real-time image generation have emerged. For instance, StreamDiffusion [28] can generate images at a speed of 100 FPS. This enables Stable Diffusion to meet the real-time requirements of video streaming.
However, due to the inability to precisely control image generation, Stable Diffusion cannot meet the fidelity requirements of video streaming. Specifically, Stable Diffusion only defines the generation process from prompt to image, without considering the inverse process of extracting prompts from images. The most straightforward solution is to automatically extract text descriptions for target images using Image/Video Captioning algorithms [46, 23] and use these extracted descriptions as prompts to generate images. Although the generated images can semantically align with the target images, these extracted descriptions are always too high-level, resulting in generated images lacking many details or having large structural differences with the target images. In fact, even if human intelligence is used to manually describe images, it is impossible to generate images that are pixel-aligned with the target images, as shown in Figure 1. In addition to using text as prompts, ControlNet [54] can utilize images (such as masks) as prompts, controlling Stable Diffusion to generate images that align with the contours of the prompt image. However, apart from the contours, details such as color and texture cannot be aligned. In conclusion, the frames generated by Stable Diffusion cannot faithfully reproduce the original frames at the pixel level, making them unsuitable for video streaming.
To materialize the above potential benefits, Promptus is the first to successfully invert frames into prompts while ensuring pixel-level alignment. Promptus is related to some work on Text Inversion [21, 39, 25], both of which learn prompts from images in an inverse manner. However, Text Inversion aims to learn new words that represent the appearance of specific objects from images, aligning only at the semantic level. Text Inversion has also been used for image compression [35]. But similarly, the inverse prompt is only used for semantic alignment, while pixel alignment is achieved by using low-resolution images as conditions.
3 Video to Prompt Inversion
Figure 3: Workflow of Promptus’s video to prompt inversion.
This section describes how Promptus inverts frames into prompts. The workflow is illustrated in Figure 3. First, to ensure pixel alignment of the generated frames, a gradient descent-based prompt fitting framework is proposed (§3.1). Second, to control the bitrate of each prompt, a low-rank decomposition-based compression algorithm is introduced (§3.2). Third, to perform inter-frame compression on prompts, a temporally smooth prompt space is proposed (§3.3).
3.1 Gradient Descent based Prompt Fitting
Gradient Descent Framework. To obtain the inverse prompt, the most straightforward approach is to train a neural network to map the target image to a prompt. However, on one hand, this approach makes it difficult to ensure pixel alignment (like the aforementioned Captioning algorithm [46, 23]). On the other hand, as the inverse process of Stable Diffusion, this neural network needs to learn comparable knowledge, but this is very expensive (e.g., training a Stable Diffusion will cost between 600,000 and 10 million US dollars [12]). Therefore, instead of training a new neural network, we propose fully leveraging Stable Diffusion’s knowledge to infer the prompt. To this end, we adopt gradient descent to iteratively fit the prompt, with the framework shown in Figure 3. Specifically, at the beginning, the prompt is randomly initialized. Then, Stable Diffusion generates a frame based on this prompt. Since the prompt is random, the generated frame is meaningless. Third, the pixel-wise difference between the generated frame and the target frame is calculated as the loss value. Fourth, backpropagation is used to compute the gradient of the loss value with respect to the prompt. Finally, the prompt is updated using gradient descent. The above steps are iteratively executed until the loss value is sufficiently small, and the resulting prompt can satisfy pixel-aligned generation. In the above steps, Stable Diffusion is pre-trained and frozen, so it has prior knowledge. This knowledge is gradually distilled into the prompt through gradient descent fitting.
To realize the aforementioned framework, there are several key components:
Single-step denoising to avoid higher-order derivatives. As described in §2.2, Stable Diffusion uses iterative denoising to generate images. Therefore, the prompt recursively affects the generated image. This causes the gradient of the loss value with respect to the prompt to involve the computation of higher-order derivatives (such as 20th order), which introduces prohibitive computational and memory overhead. Therefore, instead of using the traditional Stable Diffusion, we adopt SD Turbo [9], a variant that can generate frames through single-step denoising. The adoption of SD Turbo allows gradient descent to only compute the first-order derivatives of the prompt, greatly improving efficiency. Although the quality of the generated frames is slightly weaker than the traditional iterative denoising Stable Diffusion, these quality losses can be compensated for in an end-to-end manner during prompt fitting.
Employing embeddings as prompts to avoid non-differentiability. Computing the gradient of the prompt through backpropagation requires the forward computation from the prompt to the loss value to be completely differentiable. However, as shown in Figure 2, the forward computation includes the CLIP module [36], which converts text from discrete natural language to continuous embeddings. This step involves indexing and table lookup, making it non-differentiable. Gradients cannot be propagated to the text, preventing gradient descent. To address this, we discard the non-differentiable CLIP module and directly use embeddings as prompts for conditioning Stable Diffusion. In this case, the forward computation is fully differentiable, allowing gradient descent to be performed on the embeddings. In the following sections, prompt refers to the embedding rather than the text.
Using a noisy previous frame instead of random noise to reduce latent space distance. According to §2.2, in addition to the prompt, the input random noise also affects the frames generated by Stable Diffusion. With the same prompt, different input noise results in different generated frames. At a high level, the input noise can be viewed as a point in the latent space, and the denoising process of Stable Diffusion actually moves this point under the control of the prompt. Therefore, the goal of Promptus is to find the inverse prompt that can move the point represented by the noise to the point of the target image in the latent space. Since we adopt a single-step denoising Stable Diffusion, if the random noise is far from the target image in the latent space, this movement cannot be completed in a single step, making it impossible to fit the inverse prompt. As shown in Figure 4(a), using random noise as input, after the loss value converges, the generated frame still has noticeable differences from the target frame, including blurring, noise artifacts, and inconsistent details. Therefore, we need to reduce the distance between the input noise and the target image in the latent space. We observe that in a video, adjacent frames are close in the latent space. Thus, we manually add noise to the previous frame as follows:
$N^{t}=(1-\gamma)*Z^{t-1}+\gamma*N^{0}$ (1)
Where $Z^{t-1}$ is the previous frame in the latent space. $N^{0}$ is a fixed noise. $\gamma$ is a hyperparameter that controls the degree of noise addition, which we set to $0.95$ in the experiment. $N^{t}$ is used as the noise input to Stable Diffusion for denoising the current frame. As shown in Figure 4(c), compared to random noise, the noisy previous frame can reduce the distance in the latent space, resulting in a generated frame that better matches the target frame.
Combining reconstruction and perceptual loss functions. To achieve pixel-aligned supervision, the most intuitive loss function is the per-pixel reconstruction loss, such as MSE. These reconstruction losses attempt to minimize the error of each pixel, while errors in high-frequency details and edges often lead to large pixel errors. Therefore, to reduce the overall error, reconstruction losses tend to abandon the fitting of edges and details, resulting in overly smooth and blurry images, as shown in Figure 4(b). To make the generated frames sharp and clear, one approach is to use perceptual loss instead of reconstruction loss, such as LPIPS [55]. Perceptual loss is based on deep learning and can estimate the subjective quality of images as perceived by the human eye. Since the human eye is highly sensitive to image details, perceptual loss can make the generated images sharper with richer details. However, perceptual loss aims to maximize the overall subjective quality of the image without focusing on the exact consistency of each pixel, leading to misalignment between the generated and target images. Therefore, to simultaneously ensure pixel alignment and subjective quality, we combine the reconstruction and perceptual loss as the fitting loss $D$:
$D=\alpha*D_{rec}(x,x_{gt})+(1-\alpha)*D_{per}(x,x_{gt})$ (2)
where $D_{rec}$ represents the reconstruction loss, which is MSE by default. $D_{per}$ represents the perceptual loss, which is LPIPS by default. $\alpha$ is a hyperparameter that balances pixel alignment and perceptual quality. In our experiments, we set $\alpha$ to 0.8. The final result is shown in Figure 4(c). It can be observed that by jointly optimizing the reconstruction loss and the perceptual loss, the images generated by Promptus ensure pixel-level alignment while maintaining a sharp appearance, making them almost identical to the ground truth images.
Figure 4: Visualization of prompt fitting results. (a) Using random noise as input. (b) Only using MSE as the loss function. (c) Ours. (d) Ground Truth. The results demonstrate that the noisy previous frame and the perceptual loss both contribute to the visual quality.
3.2 Low-rank Decomposition based Prompt Bitrate Control
According to §3.1, Promptus uses embeddings as prompts instead of text. However, embeddings are $m*n$ matrices (e.g., $1024*77$), where each element is a high-bit floating-point number (e.g., 32-bit float type), resulting in a much higher bitrate. To reduce the bitrate of prompts, Promptus has two directions: First, dimensionality reduction decreases the number of parameters in the prompt. Second, quantization reduces the number of bits for each parameter in the prompt.
Low-rank matrix decomposition. To perform dimensionality reduction on the prompt, the most straightforward method is to first fit the complete prompt and then perform dimensionality reduction algorithms such as SVD (Singular Value Decomposition) or PCA (Principal Component Analysis) on it. However, these methods only perform dimensionality reduction based on the data distribution of the prompt, ignoring the impact of prompt degradation on the generation results of Stable Diffusion. This inevitably leads to a degradation in the quality of the generated images. Therefore, instead of performing explicit dimensionality reduction, Promptus proposes to directly fit a low-dimensional prompt end-to-end, thereby reducing the quality degradation caused by dimensionality reduction. To achieve this, Promptus integrates the inverse process of CANDECOMP/PARAFAC decomposition [22] into gradient descent fitting. Specifically, Promptus calculates the embedding $c$ as follows:
$\vspace{-6mm}c=\frac{u*v}{\sqrt{r}}$ (3)
where $u$ and $v$ are two low-rank factor matrices, and $r$ is the rank of the embedding. $u$ and $v$ compose the embedding $c$ through outer product and normalization. At this point, the embedding $c$, as an intermediate variable, is no longer fitted or stored. $u$ and $v$, as the new representation of the prompt, will be randomly initialized and fitted.
The rank $r$ determines the trade-off between bitrate and quality. Compared to the embedding size of $m*n$ (e.g., $1024*77$), the total size of $u$ and $v$ is $(m+n)*r$. Therefore, reducing $r$ significantly lowers the bitrate. However, on the other hand, when Rank $r$ is smaller, the embedding is constrained to be a low-rank matrix, resulting in weaker representational capability and inability to fit high-frequency details in the image, as shown in Figure 7. Consequently, it is necessary to dynamically select the most appropriate $r$ based on the currently available network bandwidth to trade off between bitrate and quality.
Fitting-aware quantization. Although the number of parameters in the prompt has been significantly reduced through low-rank matrix decomposition, each parameter in $u$ and $v$ is still a high-bit floating-point number (such as 32-bit float type). Therefore, to further reduce the bitrate, it is necessary to quantize $u$ and $v$, reducing the number of bits for each parameter (such as lowering it to 8 bits). The most straightforward approach is to first fit $u$ and $v$ and then perform quantization. However, this post-quantization technique inevitably leads to quality loss. To address this, Promptus incorporates quantization into the fitting process, automatically compensating for the quantization loss through end-to-end gradient descent. We tested the impact of different quantization configurations on quality. Compared to traditional post-quantization, our method can reduce the number of bits from 32 to 8 with almost no quality loss.
3.3 Prompt inter-frame compression based on temporal smoothing
Figure 5: At the receiver, Promptus approximates the unreceived frames by linear interpolation in the prompt space.
According to §3.1, Promptus fits each frame of the video as an independent prompt, without considering the correlation between prompts across frames. Therefore, during video streaming, Promptus needs to transmit a prompt for each frame, resulting in a linear increase in bitrate as the frame rate rises. However, codec-based streaming can avoid this problem through inter-frame compression. Is it possible to perform inter-frame compression on prompts as well? The most straightforward solution is to reshape the prompt of each frame into a two-dimensional matrix and encode it using a video codec. However, unlike conventional videos, we found that this reshaped prompt looks like random noise, lacking any patterns, structures, or smooth regions, causing video codecs to no longer work.
For inter-frame compression of prompts, our insight is: prompts are high-level semantics, so the prompts of continuous video frames should change continuously. If two temporally close frames are also sufficiently close in the prompt space, then the prompts of the frames between these two frames can be approximated by linear interpolation. With this, during streaming, we only need to sparsely transmit the prompts of a few frames (as keyframes), and the prompts of the remaining frames can be obtained by linear interpolation of the keyframe prompts, as illustrated in Figure 5. Since only a small portion of keyframes need to transmit prompts, inter-frame compression is achieved.
To ensure that adjacent frames are sufficiently close in the prompt space, we add temporal smoothing regularization to the embedding during fitting, as follows:
$\lambda=\left\|c^{t}-c^{t-1}\right\|_{2}$ (4)
Here, $c^{t}$ represents the embedding of the frame currently being fitted, while $c^{t-1}$ denotes the embedding of the previously fitted frame. The final loss function $L$ is as follows:
$L=\beta*D+(1-\beta)*\lambda$ (5)
Where $\beta$ is a hyperparameter used to balance the fitting loss and the temporal smoothing regularization. In our experiments, we set it to 0.2.
Figure 6: Visualization of the prompt interpolation results. When applying temporal smoothing regularization, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video.
Temporal smoothing regularization works, as shown in Figure 6. Without temporal smoothing regularization, the interpolation results suffer from severe distortions, disrupting the video’s content and structure. With temporal smoothing regularization, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video.
4 Prompt Streaming System Design and Implementation
§3 shows how Promptus inverts videos into pixel-aligned prompts. In this section, we will further describe how Promptus utilizes this Prompt Inversion technique to design and implement a streaming system. Specifically, it will be divided into two parts: the sender side and the receiver side.
4.1 The sender side
On the sender side, Prompt Inversion replaces video encoding. At a high level, Promptus performs Prompt Inversion on raw frames as target images. The obtained prompts are then streamed to the receiver side for image generation and playback. The details are as follows:
Initialization for the first frame. As stated in §3.1, to improve the quality of the generated images, Promptus uses the noisy previous frame as the input noise for Stable Diffusion. However, for the first frame of the video (the frame index starts from 1), there is no previous frame $Z^{0}$. To address this, Promptus uses the VAE’s Encoder to map the first frame itself to the latent space, obtaining $Z^{0}$. Then, noise is added to $Z^{0}$ according to Equation 1. The noisy $N^{1}$ will be used as the input noise for Prompt Inversion of the first frame. Since the purpose of using the noisy previous frame is to reduce the distance between the input noise and the target image in the latent space, the noisy current frame naturally works as well. Note that this noise $Z^{0}$ will also be sent to the receiver side along with the prompt for generating the first frame image. On the other hand, due to the absence of a previous frame, Promptus does not apply temporal smoothing regularization for the Fitting of the first frame.
Re-initialization for abrupt scene changes. When the video undergoes drastic changes, such as suddenly switching to a new scene, the content difference between the previous frame and the current frame becomes significant, and the distance in the latent space is no longer close. In this case, using the noisy previous frame as described in §3.1 will no longer work. Therefore, Promptus detects abrupt scene changes. Specifically, Promptus calculates the distance between the current frame and the previous frame in the latent space, and when this distance exceeds a threshold, it is considered an abrupt scene change. In this situation, Promptus treats the video after the scene change as a new video and performs the aforementioned first frame initialization again.
Sparse prompt streaming. As stated in §3.3, through temporal smoothing regularization, the prompts of most frames can be approximated by linear interpolation of the prompts of keyframes. Therefore, Promptus defines the keyframe interval as $K$ (the choice of $K$ is discussed in §5.2), adding one keyframe every $K$ frames. Only the prompts of keyframes will be streamed to the receiver. Note that when the aforementioned abrupt scene changes occur, the difference between frames in the latent space is significant, and the interpolation of prompts no longer works. In this case, the last frame before the scene change will also become a keyframe and be sent. The frames after the scene change will start a new count.
Low-rank based adaptive bitrate. According to §3.2, the higher the bitrate of the prompt, the higher the quality of the generated image. Therefore, in prompt streaming, it is necessary to increase the prompt bitrate as much as possible while avoiding network congestion to maximize the user experience. Promptus adopts the WebRTC [1] framework for prompt streaming, which dynamically probes the currently available bandwidth. Consequently, Promptus adaptively adjusts the rank to make the prompt bitrate close to the probed bandwidth. It is worth noting that the prompts transmitted by Promptus are not encoded. Therefore, compared to codec-based streaming, Promptus can precisely control the bitrate.
4.2 The receiver side
On the receiver side, Stable Diffusion’s prompt-to-image generation replaces video decoding. The details are as follows:
Video generation based on sparse prompts. After receiving the prompt of keyframe $i$, the receiver first performs linear interpolation between the prompts of adjacent keyframes $i-k$ and $i$ to approximate the prompts of the intermediate $K-1$ frames. Afterward, these prompts are used for Stable Diffusion’s image generation. As described in §3.1, the generated $Z$ for each frame will be noised and used as the input noise for the generation of the next frame. Note that since the ratio of the noise (95%) in the noisy previous frame is much larger than the ratio of the previous frame (5%), the error or complete loss of the previous frame (earlier frames can be used) has almost no impact on the generation of subsequent frames.
Real-time video generation. The frame rate of video playback at the receiver depends on Promptus’s image generation speed, making real-time generation crucial. Promptus’s image generation consists of five parts: prompt dequantization, prompt composition, prompt interpolation, adding noise to previous frame and Stable Diffusion image generation. The first four parts only involve simple linear calculations, so their time consumption can be ignored. The speed of Stable Diffusion becomes the bottleneck. To address this, we adopt the StreamDiffusion [28], which accelerates Stable Diffusion image generation to 100 FPS through batch processing. By this, Promptus achieves real-time video generation on the receiver.
5 Evaluation
Our four main evaluation results are as follows: First, Promptus achieves scalable bitrate, and its quality advantage over the baselines becomes more significant as the target bitrate decreases. Second, Promptus is general to different video domains, and the more complex the video content, the greater the quality advantage of Promptus compared to the baselines. Third, under real-world network traces, Promptus enhances the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Fourth, Promptus can generate videos from prompts in real-time at a speed exceeding 150 FPS.
5.1 Experiment Setup
Test videos: To validate the generalizability of Promptus across different video domains, we selected 7 videos from 4 datasets with vastly different content, as summarized in Table 1. Specifically, the domains of these videos span natural landscapes and human activities, outdoor long-range scenes and indoor close-up scenes, real-world scenes and CG-synthesized scenes, 3D gaming scenes and 2D animations. All videos are cropped and resized to a resolution of 512*512, with a frame rate of 30 FPS. Since Promptus uses a pre-trained and frozen Stable Diffusion, it does not involve model training and does not require training videos. However, for fairness, we still checked that the above test videos were not used during the pre-training of Stable Diffusion.
Table 1: Test videos summary
Dataset #Videos #frames Description
QST [53] 2 300 Natural landscapes, outdoor distant view
UVG [32] 2 300 Human activity, face, hand, indoor close-up
GTA-IM [17] 2 300 3D Game recording, CG-synthesized scenes
Animerun [43] 1 60 2D animation, cartoon
Total 7 960
Baselines: We compare Promptus with two baselines: H.265 [5] and VAE [27, 38]. H.265 is an advanced traditional codec that achieves compression by utilizing hand-designed intra-frame prediction, motion compensation, and transform coding techniques. Moreover, H.265 can control the bitrate of the encoded video by adjusting the quantization parameters, balancing quality and bitrate. Since it does not involve learnable parameters, it has good generality. VAE is a deep learning-based neural codec. Both its encoder and decoder are trainable neural networks. The encoder maps the input image to a low-dimensional latent variable, while the decoder reconstructs the image from the latent variable. The dimensionality of the latent variable is usually much smaller than that of the original image, thus achieving compression. For inter-frame compression of VAE, we first quantize the latent variables of each frame and then encode them into a video using H.265. The quantization process causes almost no quality degradation, and the quality degradation is mainly due to video encoding. Therefore, we adjust the target bitrate of video encoding to balance the quality and bitrate of VAE. For fairness, the training set of VAE is the same as that of Stable Diffusion, both consisting of 5.85 billion images [6].
Metric: To evaluate video quality, we adopt the LPIPS [55] (Learned Perceptual Image Patch Similarity) instead of the traditional SSIM and PSNR. This is because LPIPS has a significantly higher correlation with human subjective ratings compared to SSIM and PSNR [51, 55], which means that LPIPS better reflects human subjective perception of video quality. LPIPS uses a pre-trained neural network to extract layer-wise features from the original image and the distorted image, and then calculates the normalized L2 distance between the features as the LPIPS value. Therefore, a smaller LPIPS value indicates a higher quality of the distorted image. Specifically, we use VGG [41] pre-trained on ImageNet [18] as the feature extractor for LPIPS.
Network trace: We collected 6 network traces from real-world scenarios such as subways, driving, and walking, under 2G, 3G, and 4G networks. The traces consisted of 1 from a 4G network, 2 from 3G, and 3 from 2G. Each trace lasted between 5 and 30 seconds, totaling over 100 seconds. Our traces aimed to test various weak network conditions such as poor signal coverage, network overload, high-speed movement, and frequent switching of mobile communication networks. The average bandwidth per second ranged between 50 kbps and 4000 kbps, while the one-way network delay was approximately 30 ms to 100 ms.
Testbed: To test the streaming performance, we use WebRTC [1] to deploy Promptus and the baselines. WebRTC is a real-time video communication framework that dynamically probes the currently available bandwidth to avoid congestion. To simulate real network conditions, we use Mahimahi [34] to replay the above network traces. The queue length of Mahimahi is set to 60, and the drop-tail strategy is adopted.
5.2 Trade off between bitrate and quality
This section demonstrates how different parameter configurations affect the quality-bitrate tradeoff of Promptus.
Figure 7: Visualization of the fitting results at different ranks. It can be seen that as the rank increases, the prompt can fit more details. When the rank is 4, the earrings are lost. When the rank increases to 16, the earrings are successfully fitted.
Prompt rank. We illustrate the variation in visual quality of Promptus under different ranks, as depicted in Figure 8. It indicates that the higher the prompt rank, the higher the quality of Promptus. For example, when the prompt rank increases from 4 to 16, the LPIPS decreases from 0.265 to 0.221 (on the line with a keyframe interval of 1). This is because the larger the rank, the stronger the representational capability of the prompt, allowing it to fit the target image more accurately, as described in §3.2. We present a visualization of the fitting results at different ranks, as shown in Figure 7. It can be seen that as the rank increases, the prompt can fit more details. For instance, when the rank is 4, the earrings are lost in the fitting result, while when the rank increases to 16, the earrings are successfully fitted.
Keyframe interval. Figure 8 also shows the impact of different keyframe intervals on quality. The smaller the keyframe interval, the higher the quality of Promptus. For example, when the prompt rank is 16, reducing the keyframe interval from 10 to 1 decreases the LPIPS from 0.274 to 0.221. This is because, when the interval is smaller, the distance between keyframes in the prompt space is smaller. Since we constrain the prompt space to be temporally smooth, the linear interpolation of keyframe prompts can more accurately approximate the prompts of intermediate frames, as described in §3.3.
Figure 8: The impact of prompt rank and keyframe interval on visual quality. It indicates that visual quality improves with increasing prompt rank and decreasing keyframe interval.
Quality-bitrate tradeoff. Increasing the prompt rank and reducing the keyframe interval lead to a rapid rise in bitrate while improving quality. Therefore, we present the tradeoff between bitrate and quality for Promptus, as shown in Figure 9. First, it illustrates that overall, the quality monotonically increases with the increase in bitrate. Thus, to optimize quality, Promptus sends prompts at a bitrate closest to the available bandwidth. Second, Promptus can achieve scalable bitrates. For example, by adjusting the rank in the range of 4 to 32 and the keyframe interval in the range of 2 to 8, Promptus’s bitrate spans from 113 kbps to 4284 kbps. Third, at the same bitrate, the quality varies for different configurations. For instance, when the bitrate is 550 kbps, the configuration with a rank of 8 and a keyframe interval of 4 has an LPIPS of 0.255, while the configuration with a rank of 16 and a keyframe interval of 8 has an LPIPS of 0.264, which is lower in quality than the former. So at the same bitrate, Promptus tends to choose configurations with smaller keyframe intervals for streaming.
Figure 9: The tradeoff between bitrate and quality for Promptus. The quality monotonically increases with the increase in bitrate. Besides, Promptus can achieve scalable bitrates.
5.3 Compression efficiency
This section demonstrates the compression efficiency of Promptus. Figure 10 shows the CDF of the frame quality for Promptus and baselines under 4 bitrate levels. Since a lower LPIPS represents better visual quality, a leftward shift of the curve in the figure represents more high-quality frames, and thus higher compression efficiency. We also calculate the average LPIPS for each method, represented by the vertical lines in the figure.
(a) 140 kbps
(b) 280 kbps
(c) 360 kbps
(d) 540 kbps
Figure 10: Frame quality CDF and mean of Promptus and baselines at four target bitrates. It indicates Promptus achieves better compression efficiency across all bitrate levels. Besides, the lower the bitrate, the greater the advantage of Promptus.
First, Promptus achieves better compression efficiency across all bitrate levels. For example, in Figure 10, the curves of Promptus are all to the left of the baseline curves. Second, the lower the bitrate, the greater the advantage of Promptus. At a high level, as the bitrate decreases from 540 kbps to 140 kbps, the distance between Promptus’s curves and the baselines’ curves gradually widens. At a low level, when the bitrate is 540 kbps, the average LPIPS of Promptus is 0.018 and 0.085 lower than VAE and H.265, respectively. When the bitrate is 140 kbps, this difference further increases to 0.139 and 0.118. This is because when the bitrate is reduced, H.265 uses coarser quantization and loses many high-frequency details, resulting in blurriness and block artifacts in the video, which significantly impairs the perceptual quality. VAE mitigates this phenomenon by mapping images to a low-dimensional latent space, reserving more bitrate for video coding. Therefore, at most bitrates, the quality of VAE is superior to H.265. However, at extremely low bitrates (such as 140 kbps), the latent space inevitably introduces distortions caused by video coding. At this point, the VAE Decoder introduces a large number of errors when reconstructing the images, causing a significant decrease in VAE’s quality. On the other hand, when the bitrate is reduced, Promptus reduces the representational capability of the prompt rather than degrading the video quality. This prevents Promptus from accurately describing the video content, resulting in slight misalignments in the generated frames. However, thanks to the inherent image generation capability of Stable Diffusion, Promptus’s frames still have good sharpness and details, and thus the perceptual quality is better than VAE and H.265.
5.4 Generality
This section demonstrates the generality of Promptus across different domains. Figure 12 shows the mean of the frame quality for Promptus and two baselines on four different datasets (as described in Table 1).
Figure 11: Visualization of the compression results on different datasets. It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
Figure 12: Mean frame quality of Promptus and baselines on 4 different datasets. It demonstrates the generality of Promptus across different domains. Besides, the more high-frequency details a video has, the greater the advantage of Promptus.
First, Promptus achieves better compression efficiency on each dataset. This is because, as shown in Figure 12, Promptus achieves lower average LPIPS compared to the baselines on each dataset. To intuitively demonstrate this improvement, we also visualize the compression results of the three methods on the four datasets at lower bitrates (such as 225 kbps), as shown in Figure 11. It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
Second, the more high-frequency details a video has, the greater the advantage of Promptus. For example, for the Animerun dataset with fewer details, the LPIPS of Promptus is 0.015 lower than H.265, which is not a significant advantage. However, for the detail-rich UVG, this difference further expands to 0.121. This is because for 2D animations with large areas of solid colors and simple details, H.265’s intra-frame prediction, block partitioning, and motion compensation techniques can handle them well, so Promptus’s performance gain is small. For detail-rich real-world videos, H.265 discards more high-frequency information during compression, thus damaging the perceptual quality (especially at lower bitrates). Although Promptus also loses high-frequency information from the original image, Stable Diffusion completes the lost high-frequency information based on prior knowledge during generation, resulting in a smaller perceptual quality loss.
5.5 Performance on real-world traces
This section demonstrates the performance of Promptus under real network traces. We use the traces and testbed described in §5.1 to run Promptus and two baselines. Figure 13 shows the CDF and mean of the frame quality. Since the total length of the traces is greater than the total length of the test videos, we loop the test videos to run through the entire traces.
Figure 13: The performance of Promptus under real network traces. On one hand, Promptus’s quality is overall higher than the baselines. On the other hand, Promptus can significantly reduce the ratio of severely distorted frames.
First, Promptus’s quality is overall higher than the baselines. For example, the mean LPIPS of Promptus is 0.111 and 0.092 lower than VAE and H.265, respectively. Second, Promptus can significantly reduce the ratio of severely distorted frames. For instance, only 5.2% of frames in Promptus have LPIPS higher than 0.32, while VAE and H.265 have 94.5% and 96.9%, respectively. These improvements are partly due to Promptus’s excellent compression efficiency, enabling it to provide higher perceptual quality at the same bitrate. On the other hand, since Promptus sends raw prompts without encoding (such as entropy coding), it can precisely control the target, thus making full use of bandwidth.
5.6 Overhead
In this section, we analyze the overhead of Promptus’s generation. We conduct tests on an Nvidia 4090D GPU, using CUDA to accelerate Promptus. The resolution of the generated video is 512*512, and the rank of the prompt is 8.
Table 2: Overhead of generating a frame
Steps Time (ms) Memory (MB)
Prompt dequantization 0.016 -
Prompt composition 0.025 -
Prompt interpolation 0.013 -
Noised previous frame 0.012 -
Stable Diffusion generation 6.160 -
Total 6.226 8952
Table 2 shows the fine-grained overhead of each step in Promptus’s image generation. Specifically, from receiving a prompt to generating a frame, Promptus includes the following steps: prompt dequantization, prompt composition, prompt interpolation, adding noise to the previous frame, and Stable Diffusion image generation. Among them, most steps only involve simple linear computations, so the time overhead is almost negligible. In contrast, Stable Diffusion image generation accounts for the vast majority of the time overhead. Therefore, we adopt StreamDiffusion [28], which accelerates Stable Diffusion to real-time through batch processing techniques and implementation optimizations. The difference is that compared to the original StreamDiffusion, Promptus removes the CLIP module that converts text to embeddings, thus further accelerating image generation to 6.160 ms. In summary, the total time overhead of Promptus’s image generation is 6.226 ms, achieving real-time.
6 Limitation
The time overhead of prompt fitting. Although Promptus can achieve real-time video generation at the receiver, prompt fitting cannot be performed in real-time at the sender. This is because prompt fitting requires iterative gradient descent. Although a single iteration is fast (on the same order of magnitude as the time overhead of real-time generation), the total time overhead is high due to the large number of iterations required for convergence (such as 500 iterations). To address this, using more efficient gradient descent algorithms to reduce the number of iterations required for convergence is expected to accelerate prompt fitting to real-time.
The latency of prompt interpolation. At the receiver side, Promptus obtains the prompts of intermediate frames through prompt interpolation of keyframes. This means that if an intermediate frame needs to be generated and played, it is necessary to wait until the subsequent keyframe is received, which introduces additional latency. To address this issue, designing keyframe extrapolation algorithms to replace interpolation is a future research direction.
Non-uniform keyframes. According to §4.1, Promptus sends keyframes uniformly based on the keyframe interval. However, different segments of a video often have different rates of change. For rapidly changing segments, the distance between keyframes in the prompt space is large, resulting in a poor approximation of intermediate frames using linear interpolation. A solution is to reduce the keyframe interval and send keyframes more densely. However, there also exist smoothly changing segments in the video, where densely sending keyframes brings little improvement to quality, thus wasting bandwidth. In summary, sending keyframes uniformly in this paper is inefficient. In the future, adaptive keyframe needs to be designed to transmit densely in rapidly changing segments and sparsely in smoothly changing segments.
7 Conclusion
In this paper, we propose Promptus, a novel system that replaces video streaming with prompt streaming by inverting video frames into prompts for Stable Diffusion. To ensure pixel alignment, a gradient descent-based prompt fitting framework is proposed. To achieve adaptive bitrate for prompts, a low-rank decomposition-based bitrate control algorithm is introduced. For inter-frame compression of prompts, a temporal smoothing-based prompt interpolation algorithm is proposed. Evaluations across various video domains and real network traces demonstrate Promptus can enhance the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Moreover, Promptus achieves real-time video generation from prompts at over 150 FPS. By pioneering the replacement of video codecs with prompt inversion and introducing prompt streaming, Promptus revolutionizes efficient video communication, surpassing the limitations of traditional approaches and paving the way for a new era of video streaming.

### v2

@@
 With the exponential growth of video traffic, traditional video streaming systems are approaching their limits in compression efficiency and communication capacity.
-To further reduce bitrate while maintaining quality, we propose Promptus, a disruptive novel system that streaming prompts instead of video content with Stable Diffusion, which converts video frames into a series of "prompts" for delivery.
-To ensure pixel alignment, a gradient descent-based prompt fitting framework is proposed.
-To achieve adaptive bitrate for prompts, a low-rank decomposition-based bitrate control algorithm is introduced.
-For inter-frame compression of prompts, a temporal smoothing-based prompt interpolation algorithm is proposed.
-Evaluations across various video domains and real network traces demonstrate Promptus can enhance the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%.
-Moreover, Promptus achieves real-time video generation from prompts at over 150 FPS.
-To the best of our knowledge, Promptus is the first attempt to replace video codecs with prompt inversion and the first to use prompt streaming instead of video streaming.
-Our work opens up a new paradigm for efficient video communication beyond the Shannon limit.
+To further reduce bitrate while maintaining quality, we propose Promptus, a disruptive semantic communication system that streaming prompts instead of video content, which represents real-world video frames with a series of "prompts" for delivery and employs Stable Diffusion to generate videos at the receiver.
+To ensure that the prompt representation is pixel-aligned with the original video, a gradient descent-based prompt fitting framework is proposed.
+Further, a low-rank decomposition-based bitrate control algorithm is introduced to achieve adaptive bitrate.
+For inter-frame compression, an interpolation-aware fitting algorithm is proposed.
+Evaluations across various video genres demonstrate that, compared to H.265, Promptus can achieve more than a 4x bandwidth reduction while preserving the same perceptual quality.
+On the other hand, at extremely low bitrates, Promptus can enhance the perceptual quality by 0.139 and 0.118 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%.
+Our work opens up a new paradigm for efficient video communication.
+Promptus is open-sourced at: https://github.com/JiangkaiWu/Promptus.
 1 Introduction
-With the rapid development of streaming applications (such as video-on-demand [2, 8], live video [14, 7], video conferencing [15, 11], cloud gaming [3, 13], etc.), the traffic of network video has been continuously growing.
-To reduce traffic, video codecs represented by VP8 [16], VP9 [33], H.264 [4] and H.265 [5] are widely used to compress videos.
+With the rapid development of streaming applications (such as Youtube, Netflix), the traffic of network video has been continuously growing.
+To reduce traffic, video codecs represented by VP8 (bankoski2011technical, ), VP9 (mukherjee2015technical, ), H.264 (264, ) and H.265 (265, ) are widely used to compress videos.
 These codecs achieve compression by removing spatial and temporal redundancies.
-However, these redundancies are limited, so there is an upper bound on the compression ratio (subject to the Shannon limit [40]).
-In order to further compress the video, non-redundant content in the video will be discarded, which will greatly degrade the video quality, such as causing blurring and blocking artifacts.footnotetext: Promptus will be released at https://pku-netvideo.github.io/ once published.
+However, these redundancies are limited, so there is an upper bound on the compression ratio (subject to the Shannon limit (shannon1948mathematical, )).
+In order to further compress the video, non-redundant content in the video will be discarded, which will greatly degrade the video quality, such as causing blurring and blocking artifacts.
+In recent years, some deep learning-based codecs (lu2019dvc, ; lin2020m, ; djelouah2019neural, ) and streaming frameworks (zhou2022cadm, ; jiang2022wireless, ; sivaraman2024gemino, ; li2023reparo, ) have been proposed to improve compression ratio, but they are limited either by performance or by generality.
 Figure 1: Promptus can invert a given image into a prompt.
 Based on this prompt, Stable Diffusion can generate an almost identical image to the original.
-In contrast, existing methods can only generate semantically similar images.
-To improve compression efficiency, some deep learning-based codecs and streaming frameworks have been proposed in recent years.
-Neural-embedded codecs [31, 30, 19, 37, 47] replace handcrafted modules (such as motion vector estimation [45]) in traditional codec frameworks with neural networks (such as autoencoders), learning more intelligent and efficient compression capabilities through end-to-end optimization.
-However, due to still following the traditional encoding framework, the improvement in compression ratio is limited.
-Neural-enhanced streaming [50, 49, 26, 52, 56] transmits distorted low-bitrate videos and then restores high-fidelity details using post-processing algorithms such as super-resolution.
-However, these post-processing algorithms rely on prior knowledge learned from training sets.
-Due to the domain gap, their performance will degrade in unseen scenes [48].
-Generative streaming [24, 44, 42, 29] transmits small-sized semantic information, and the receiver can generate video based on these semantics.
-However, most generative algorithms are designed for specific tasks (e.g., talking video generation driven by facial keypoints [24, 44]), limiting their generality.
-With the popularity of AIGC (AI-generated content), Stable Diffusion [38, 10, 9] has attracted extensive attention thanks to its powerful text-to-image generation capability.
-By pre-training on an internet-scale dataset [6], Stable Diffusion learns prior knowledge of all human visual domains, and simultaneously learns the mapping from text to images.
+In contrast, existing methods can only generate semantically similar images, while the differences at the pixel level are substantial.
+In this way, Promptus streams prompts instead of videos, significantly reducing bandwidth overhead.
+With the popularity of generative AI, Stable Diffusion (rombach2022high, ; sd, ) has attracted extensive attention thanks to its powerful text-to-image generation capability.
+By pre-training on an internet-scale dataset (LAION, ), Stable Diffusion learns prior knowledge of nearly all human visual domains, and simultaneously learns the mapping from text to images.
 Therefore, Stable Diffusion can generate high-fidelity images based on a brief prompt composed of a few words.
@@
 During streaming, the sender streams prompts instead of streaming encoded videos and the receiver generates videos instead of decoding videos.
-In this way, the traffic of network video is reduced from the video scale to the text scale, achieving communication efficiency beyond the Shannon limit.
-In this paper, we propose Promptus, a system that inverts video frames into Stable Diffusion prompts, achieving ultra-low bitrate video streaming.
+In this way, the traffic of network video is reduced from the video scale to the text scale, greatly improving video communication efficiency.
+In this paper, we propose Promptus, a system that represents real-world video frames with Stable Diffusion prompts, achieving ultra-low bitrate video streaming.
 To bring this vision to fruition, we address the following technical challenges:
@@
 The partial derivative of the loss with respect to the prompt will be calculated and the prompt is iteratively optimized using gradient descent.
-To implement this framework, Promptus employs single-step denoising instead of iterative denoising to avoid higher-order derivatives.
+To implement this framework, Promptus employs single-step denoising instead of iterative denoising to avoid unrolled long gradient chains.
 Second, Promptus uses embeddings as prompts instead of text to avoid non-differentiability.
@@
 When the rank is higher, the representational capability of the prompt is better, and it can describe more details in the frame, but the data size is also larger.
-Therefore, Promptus adaptively selects the prompt rank based on the currently available bandwidth.
+So Promptus adaptively selects the prompt rank based on the currently available bandwidth.
 Third, how to perform inter-frame compression on prompts.
-Promptus inverts each frame into an independent prompt without considering the correlation between prompts across frames.
-Therefore, when streaming video, Promptus needs to transmit prompts for each frame, resulting in the bitrate increasing linearly with the frame rate.
-To address this, Promptus adds a temporal smoothing regularization during prompt fitting, ensuring that temporally close frames are also sufficiently close in the prompt space.
+Promptus inverts each frame into an independent prompt without considering the correlation across frames.
+Therefore, when streaming, Promptus needs to transmit prompts for each frame, resulting in the bitrate increasing linearly with the frame rate.
+To address this, Promptus adds an interpolation-aware fitting algorithm, ensuring that consecutive frames also exhibit continuity in the prompt space.
 With this, Promptus only needs to sparsely transmit prompts for a few keyframes, while the prompts for the remaining frames can be approximated through linear interpolation of the keyframe prompts.
-We evaluated Promptus on test videos from different domains and under real network traces.
+We evaluated Promptus on test videos from different domains and under real-world network traces.
 The results show that: First, Promptus achieves scalable bitrate, and its quality advantage over the baselines becomes more significant as the target bitrate decreases.
 Second, Promptus is general to different video domains, and the more complex the video content, the greater the quality advantage of Promptus compared to the baselines.
-Third, under real-world network traces, Promptus enhances the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE [27, 38] and H.265 [5], respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%.
-Fourth, Promptus can generate videos from prompts in real-time at a speed exceeding 150 FPS.
-The contributions of this paper are summarized as follows: (1) We propose Promptus, which, to the best of our knowledge, is the first attempt to replace video codecs with prompt inversion and also the first to use prompt streaming to replace video streaming (§3).
-(2) We propose a gradient descent-based prompt fitting framework, achieving pixel-aligned prompt inversion for the first time (§3.1).
-(3) We build a video streaming system based on Promptus (§4).
-2 Motivation and related work
+Third, compared to H.265 (265, ), Promptus provides more than a 4x bandwidth reduction while preserving the same perceptual quality.
+On the other hand, at extremely low bitrates, Promptus can enhance the perceptual quality by 0.139 and 0.118 (in LPIPS (zhang2018unreasonable, )) compared to VAE (kingma2014auto, ; rombach2022high, ) and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%.
+The contributions of this paper are summarized as follows: (1) We propose Promptus, which, to the best of our knowledge, is the first attempt to replace video codecs with prompt inversion and also the first to use prompt streaming to replace video streaming.
+(2) We propose a gradient descent-based prompt fitting framework, achieving pixel-aligned prompt inversion.
+(3) We build a video streaming system based on prompts, significantly boosting video communication efficiency.
+2 Related work and Background
 2.1 Video codec and streaming
-Traditional codecs.
-In network video traffic, most of the content is encoded using traditional codecs represented by VP8 [16], VP9 [33], H.264 [4] and H.265 [5].
+Video codecs.
+In network video traffic, most of the content is encoded using traditional codecs represented by VP8 (bankoski2011technical, ), VP9 (mukherjee2015technical, ), H.264 (264, ) and H.265 (265, ).
 These traditional video codecs achieve compression primarily by eliminating redundancy in the video.
@@
 Besides, considering the similarity between consecutive video frames, the codecs employ inter-frame prediction techniques to eliminate temporal redundancy.
-For moving objects, the codecs perform motion estimation and compensation to further reduce redundancy.
 Although these techniques achieve efficient compression, the compression ratio has an upper limit due to the finite amount of redundancy.
-Further compression inevitably requires discarding some non-redundant information, resulting in a drastic decrease in video quality.
-For example, as shown in Figure 11, when the bitrate is extremely low, videos compressed using H.265 exhibit significant blurriness.
-Neural-embedded codecs.
-With the development of deep learning, handcrafted modules in traditional codecs are being replaced by neural networks [31, 30, 19, 37, 47].
+With the development of deep learning, handcrafted modules in traditional codecs are being replaced by neural networks (lu2019dvc, ; lin2020m, ; djelouah2019neural, ; rippel2019learned, ; li2024neural, ; sheng2024prediction, ).
 Compared to handcrafted modules, these embedded neural networks can learn more complex and nonlinear motion patterns and intra-frame correlations, thereby achieving lower distortion at the same bitrate.
-Furthermore, due to the ability to optimize end-to-end, neural-embedded codecs can be trained for tasks beyond compression, such as loss resilience.
-However, since neural-embedded codecs still adhere to the traditional coding framework, aiming to fully remove intra-frame or inter-frame redundancy, the improvement in compression ratio is limited.
+However, neural-embedded codecs still adhere to the traditional coding framework (aiming to remove redundancy), so the improvement is limited.
 Neural-enhanced streaming.
-In contrast to reducing redundancy, neural-enhanced streaming [50, 49, 26, 52, 56] actively discards most of the information (including non-redundant information) and transmits highly distorted low-bitrate videos.
-Then, at the receiver side, neural network-based post-processing algorithms (such as super-resolution) are used to restore high-fidelity details.
-Neural-enhanced streaming leverages the powerful image restoration and enhancement capabilities of neural networks, effectively recovering lost textures and details, thus ensuring video quality while achieving substantial compression.
-However, these post-processing algorithms rely on prior knowledge learned from training datasets, such as texture features and edge structures.
+In contrast to reducing redundancy, neural-enhanced streaming (park2023omnilive, ; zhou2022cadm, ) actively discards most of the information (including non-redundant information), transmitting highly distorted low-bitrate videos.
+Then, at the receiver, neural network-based post-processing methods (such as super-resolution) are used to restore high-fidelity details.
+Thanks to the powerful image restoration capabilities of neural networks, the lost details can be effectively recovered, thus ensuring video quality while achieving substantial compression.
+However, these post-processing neural networks rely on prior knowledge learned from training datasets.
 Due to the complex and diverse nature of real-world video, there exists domain gaps with the training set.
-The performance of these algorithms often degrades in unseen scenarios [48].
+The performance often degrades in unseen scenarios (yang2021real, ).
+A feasible solution is to fine-tune the post-processing neural networks for each video segment and send the fine-tuned neural networks along with the video, like NAS (yeo2017will, ; yeo2018neural, ; yeo2022neuroscaler, ; yeo2020nemo, ; kim2020neural, ; zhang2022yuzu, ).
+However, the transmission of the neural networks inevitably increases the overall bitrate.
+Especially for shorter videos (like Tiktok), the overhead of neural networks is unacceptable.
 Generative streaming.
-Compared to using neural networks for post-processing, generative streaming directly uses neural networks to generate videos to handle higher compression ratios [24, 44, 42, 29].
-For example, in the context of video conferencing, Face-vid2vid [44, 24] extracts facial keypoints in real-time at the sender side for transmission.
+Compared to using neural networks for post-processing, generative streaming directly uses neural networks to generate videos to handle higher compression ratios (jiang2022wireless, ; wang2021one, ; sivaraman2024gemino, ; li2023reparo, ).
+For example, in the context of video conferencing, Face-vid2vid (wang2021one, ; jiang2022wireless, ) extracts facial keypoints in real-time at the sender side for transmission.
 At the receiver side, it generates dynamic facial videos using the keypoints and static facial images.
 The bitrate of facial keypoints is much lower than that of videos, thus compressing video conferences to extremely low bitrates.
-Similarly, Gemino [42] transmits video streams at extremely low resolution.
+Similarly, Gemino (sivaraman2024gemino, ) transmits video streams at extremely low resolution.
 At the receiver side, it estimates facial motion fields from the low-resolution video, then utilizes the motion fields and high-resolution facial images to generate high-resolution facial videos.
-Instead of driving static images, Reparo [29] proposes a token-based generative streaming.
-It first uses VQGAN [20] to train a codebook of facial visual features, then maps the video to latent variables using VAE [27], and finally quantizes the latent variables into tokens according to the codebook.
-Since tokens are indices of the codebook, the bitrate is extremely low.
+Instead of driving static images, Reparo (li2023reparo, ) proposes a token-based generative streaming.
+It first uses VQGAN (esser2021taming, ) to train a codebook of facial visual features, then maps the video to latent variables using VAE (kingma2014auto, ), and finally quantizes the latent variables into token indices according to the codebook.
+Since only indices need to be transmitted, the bitrate is very low.
 In general, generative streaming can greatly compress videos, but it is often designed for specific tasks (such as video conferencing) and lacks generality.
 2.2 Stable diffusion
-Stable Diffusion [38, 10, 9] is the most popular open-source text-to-image generative model.
-It learns a denoising process from 5.85 billion image-text pairs [6], enabling it to generate high-quality images by denoising the pure noise images.
+Stable Diffusion (rombach2022high, ; sd, ) is the most popular open-source text-to-image generative model.
+It learns a denoising process from 5.85 billion image-text pairs (LAION, ), enabling it to generate high-quality images by denoising the pure noise images.
 Different noise images lead to different generated images.
-Since the noise images are randomly sampled from Gaussian distribution, the images generated by Stable Diffusion are random and uncontrollable.
-Therefore, to control the content of the generated images, Stable Diffusion also receives user-input prompts (in natural language) as the condition for denoising, thereby generating images that align with the semantic descriptions of the prompts.
-Specifically, as shown in Figure 2, $p$ represents the user-input prompt.
-Stable Diffusion first performs word embedding and semantic extraction using the CLIP model [36], converting discrete natural language into a continuous text embedding $c$ (an $m*n$ matrix).
-The text embeddings and the randomly sampled noise image $N$ are input into the denoising process of Stable Diffusion, generating the denoised $N^{\prime}$.
+Since the noise images are randomly sampled from Gaussian distribution, the generated images are random and uncontrollable.
+Thus, to control the content of the generated images, Stable Diffusion also receives user-input prompts (in natural language) as the condition for denoising, thereby generating images that align with the semantic descriptions of the prompts.
+Specifically, as shown in Figure 2, let $p$ represent the user-input prompt.
+Stable Diffusion first performs word embedding and semantic extraction using the CLIP model (radford2021learning, ), converting discrete natural language $p$ into a continuous text embedding $c$ (an $m*n$ matrix).
+The text embeddings $c$ and the randomly sampled noise image $N$ are input into the denoising process of Stable Diffusion, generating the denoised $N^{\prime}$.
 Next, $N^{\prime}$ is input into the denoising process again for $T$ iterations to obtain the denoised $Z$.
-Since the denoising process occurs in the latent space, $Z$ needs to be input into the VAE Decoder [27] to generate the image $x$ in the pixel space.
+Since the denoising process occurs in the latent space, $Z$ needs to be input into the VAE Decoder (kingma2014auto, ) to generate the image $x$ in the pixel space.
 Figure 2: How does Stable Diffusion generate high-quality images from text prompts. The details are elaborated in §2.2.
@@
 In codec-based streaming, such high-quality images require high bitrates.
-Instead of transmitting image, if we solely transmit prompts and generate images at the receiving end, the bitrate can be reduced from the image scale to the prompt scale, greatly improving communication efficiency.
+Instead of transmitting images, if we solely transmit prompts and generate images at the receiving end, the bitrate can be reduced from the image scale to the prompt scale, greatly improving communication efficiency.
+  • Stable Diffusion decouples the prompt size from the video resolution.
+A key insight in Stable Diffusion’s image generation is that the resolution of the generated images is independent of the prompt size.
+As illustrated in Figure 2, since the denoising process does not change the resolution, the resolution of the generated image $x$ is solely dependent on the resolution of the input noise image $N$.
+Since the noise image $N$ is simply obtained through random sampling from a Gaussian distribution, the resolution can be set freely.
+Given the same prompt, using noise images $N$ sampled at different resolutions will generate high-quality images that are semantically similar but differ in resolution.
+In contrast, encoder-based compression (such as VAE (kingma2014auto, ) and H.265) inevitably requires an increase in data size as the resolution rises.
+This is the fundamental difference between prompts and encoders and is the key to achieving ultra-low bitrate video communication.
   • Stable Diffusion is generalizable to most video domains.
-Trained on an internet-scale dataset [6] (5.85 billion images), Stable Diffusion has learned priors for nearly all video domains present on the internet.
+Trained on an internet-scale dataset LAION (5.85 billion images), Stable Diffusion has learned priors for nearly all video domains present on the internet.
 Consequently, it does not suffer from the domain gap issue and can generate images of any style and content.
@@
 With ongoing advancements in Stable Diffusion, variants capable of real-time image generation have emerged.
-For instance, StreamDiffusion [28] can generate images at a speed of 100 FPS.
-This enables Stable Diffusion to meet the real-time requirements of video streaming.
+For instance, StreamDiffusion kodaira2023streamdiffusion can generate images at a speed of 100 FPS.
+This enables Stable Diffusion to meet the real-time playback requirements at the receiver.
 However, due to the inability to precisely control image generation, Stable Diffusion cannot meet the fidelity requirements of video streaming.
-Specifically, Stable Diffusion only defines the generation process from prompt to image, without considering the inverse process of extracting prompts from images.
-The most straightforward solution is to automatically extract text descriptions for target images using Image/Video Captioning algorithms [46, 23] and use these extracted descriptions as prompts to generate images.
-Although the generated images can semantically align with the target images, these extracted descriptions are always too high-level, resulting in generated images lacking many details or having large structural differences with the target images.
+Stable Diffusion only defines the generation process from prompt to image, without considering the inverse process of extracting prompts from images.
+The most straightforward solution is to extract text descriptions for target images using Image/Video Captioning algorithms (yang2023vid2seq, ; hu2022scaling, ) and use these extracted descriptions as prompts to generate images.
+Although the generated images can semantically align with the target images, these extracted descriptions are always too high-level, resulting in generated images having large structural differences with the target images.
 In fact, even if human intelligence is used to manually describe images, it is impossible to generate images that are pixel-aligned with the target images, as shown in Figure 1.
-In addition to using text as prompts, ControlNet [54] can utilize images (such as masks) as prompts, controlling Stable Diffusion to generate images that align with the contours of the prompt image.
-However, apart from the contours, details such as color and texture cannot be aligned.
-In conclusion, the frames generated by Stable Diffusion cannot faithfully reproduce the original frames at the pixel level, making them unsuitable for video streaming.
-To materialize the above potential benefits, Promptus is the first to successfully invert frames into prompts while ensuring pixel-level alignment.
-Promptus is related to some work on Text Inversion [21, 39, 25], both of which learn prompts from images in an inverse manner.
-However, Text Inversion aims to learn new words that represent the appearance of specific objects from images, aligning only at the semantic level.
-Text Inversion has also been used for image compression [35].
+Besides text, ControlNet (zhang2023adding, ) can also use images (such as masks and edges) as prompts, controlling Stable Diffusion to generate images that align with the contours of the image prompts.
+However, apart from the contours, details such as color and texture cannot be aligned qiao2024latency .
+In conclusion, the frames generated by Stable Diffusion cannot faithfully reproduce the original ones at the pixel level, making them unsuitable for video streaming.
+To materialize the above potential benefits, Promptus is an initial effort to invert frames into prompts while ensuring pixel-level alignment.
+It is related to some works on Text Inversion (gal2022image, ; ruiz2023dreambooth, ; kawar2023imagic, ), both of which learn prompts from images in an inverse manner.
+However, Text Inversion aims to learn text embeddings that represent the appearance of specific objects from multiple images, aligning only at the semantic level.
+Text Inversion has also been used for image compression (pan2022extreme, ).
 But similarly, the inverse prompt is only used for semantic alignment, while pixel alignment is achieved by using low-resolution images as conditions.
@@
 First, to ensure pixel alignment of the generated frames, a gradient descent-based prompt fitting framework is proposed (§3.1).
-Second, to control the bitrate of each prompt, a low-rank decomposition-based compression algorithm is introduced (§3.2).
-Third, to perform inter-frame compression on prompts, a temporally smooth prompt space is proposed (§3.3).
+Second, to control the bitrate of each prompt, a low-rank decomposition-based adaptive bitrate algorithm is introduced (§3.2).
+Third, to perform inter-frame compression on prompts, an interpolation-aware fitting algorithm is proposed (§3.3).
 3.1 Gradient Descent based Prompt Fitting
 Gradient Descent Framework.
-To obtain the inverse prompt, the most straightforward approach is to train a neural network to map the target image to a prompt.
-However, on one hand, this approach makes it difficult to ensure pixel alignment (like the aforementioned Captioning algorithm [46, 23]).
-On the other hand, as the inverse process of Stable Diffusion, this neural network needs to learn comparable knowledge, but this is very expensive (e.g., training a Stable Diffusion will cost between 600,000 and 10 million US dollars [12]).
-Therefore, instead of training a new neural network, we propose fully leveraging Stable Diffusion’s knowledge to infer the prompt.
+In order to obtain the inverse prompt, the most straightforward approach is to train a neural network to map the target image to a prompt.
+However, on one hand, this approach makes it difficult to ensure pixel alignment (like the aforementioned Captioning (yang2023vid2seq, ; hu2022scaling, )).
+On the other hand, as the inverse process of Stable Diffusion, this neural network needs to learn comparable knowledge, but this is very expensive (e.g., training a Stable Diffusion will cost between 600,000 and 10 million US dollars (SDcost, )).
+Instead of training a new neural network, we propose fully leveraging Stable Diffusion’s knowledge to infer the prompt.
 To this end, we adopt gradient descent to iteratively fit the prompt, with the framework shown in Figure 3.
 Specifically, at the beginning, the prompt is randomly initialized.
-Then, Stable Diffusion generates a frame based on this prompt.
+Stable Diffusion generates a frame based on this prompt.
 Since the prompt is random, the generated frame is meaningless.
 Third, the pixel-wise difference between the generated frame and the target frame is calculated as the loss value.
-Fourth, backpropagation is used to compute the gradient of the loss value with respect to the prompt.
+Fourth, backpropagation is used to compute the gradient of the loss with respect to the prompt.
 Finally, the prompt is updated using gradient descent.
-The above steps are iteratively executed until the loss value is sufficiently small, and the resulting prompt can satisfy pixel-aligned generation.
+The above steps are iteratively executed until the loss is sufficiently small, and the resulting prompt can satisfy pixel-aligned generation.
 In the above steps, Stable Diffusion is pre-trained and frozen, so it has prior knowledge.
-This knowledge is gradually distilled into the prompt through gradient descent fitting.
+This knowledge is distilled into the prompt via gradient descent.
 To realize the aforementioned framework, there are several key components:
-Single-step denoising to avoid higher-order derivatives.
+Single-step denoising to avoid unrolled long gradient chains.
 As described in §2.2, Stable Diffusion uses iterative denoising to generate images.
-Therefore, the prompt recursively affects the generated image.
-This causes the gradient of the loss value with respect to the prompt to involve the computation of higher-order derivatives (such as 20th order), which introduces prohibitive computational and memory overhead.
-Therefore, instead of using the traditional Stable Diffusion, we adopt SD Turbo [9], a variant that can generate frames through single-step denoising.
-The adoption of SD Turbo allows gradient descent to only compute the first-order derivatives of the prompt, greatly improving efficiency.
-Although the quality of the generated frames is slightly weaker than the traditional iterative denoising Stable Diffusion, these quality losses can be compensated for in an end-to-end manner during prompt fitting.
-Employing embeddings as prompts to avoid non-differentiability.
+So the prompt recursively influences the generated image.
+This causes the gradient of the loss value with respect to the prompt to involve unrolling numerous iterations (e.g., 50 steps), where the accumulated long gradient chain introduces prohibitive computational and memory overhead.
+Thus, instead of using the traditional Stable Diffusion, we adopt SD (2.1) Turbo (sdturbo, ), a variant that can generate frames through single-step denoising.
+The adoption of SD Turbo allows gradient descent to require only a single forward-backward pass, greatly improving efficiency.
+Although the quality of the generated frames is weaker than the traditional Stable Diffusion, these quality losses can be compensated for in an end-to-end manner during fitting.
+Employing embeddings as prompts to prevent non-differentiability.
 Computing the gradient of the prompt through backpropagation requires the forward computation from the prompt to the loss value to be completely differentiable.
-However, as shown in Figure 2, the forward computation includes the CLIP module [36], which converts text from discrete natural language to continuous embeddings.
-This step involves indexing and table lookup, making it non-differentiable.
-Gradients cannot be propagated to the text, preventing gradient descent.
-To address this, we discard the non-differentiable CLIP module and directly use embeddings as prompts for conditioning Stable Diffusion.
-In this case, the forward computation is fully differentiable, allowing gradient descent to be performed on the embeddings.
-In the following sections, prompt refers to the embedding rather than the text.
+However, as described in §2.2, the forward computation includes the CLIP module (radford2021learning, ), which converts text from discrete natural language to continuous text embeddings.
+This conversion involves indexing and table lookup, making it non-differentiable.
+Gradients cannot be propagated to the text in natural language, preventing gradient descent.
+To address this, we discard the non-differentiable CLIP module and directly use text embeddings as prompts for conditioning Stable Diffusion.
+In this case, the forward computation is fully differentiable, allowing gradient descent to be performed on the text embeddings.
+In the following sections, prompt refers to the text in embedding rather than the text in natural language [^1] It is important to note that, even though it is not natural language, embeddings can still serve as prompts.
+On one hand, prompts can take multiple modalities, such as images (zhang2023adding, ; blattmann2023stable, ), audio (biner2024sonicdiffusion, ; tang2024any, ), embeddings (gal2022image, ; ruiz2023dreambooth, ; kawar2023imagic, ; pan2022extreme, ), fMRI (chen2024cinematic, ; takagi2023high, ), etc.
+On the other hand, the fundamental characteristic of prompts is that they serve as conditions to guide generation.
+The resolution of the generated videos is only dependent on the size of the random noise, unrelated to the prompt size, as described in §2.2.
+So prompts have the advantage of being decoupled from video resolution compared to encoding..
 Using a noisy previous frame instead of random noise to reduce latent space distance.
-According to §2.2, in addition to the prompt, the input random noise also affects the frames generated by Stable Diffusion.
+According to §2.2, in addition to the prompt, the input random noise also affects the generated frames.
 With the same prompt, different input noise results in different generated frames.
-At a high level, the input noise can be viewed as a point in the latent space, and the denoising process of Stable Diffusion actually moves this point under the control of the prompt.
+At a high level, the input noise can be viewed as a point in the latent space, and the denoising process actually moves this point under the control of the prompt.
 Therefore, the goal of Promptus is to find the inverse prompt that can move the point represented by the noise to the point of the target image in the latent space.
-Since we adopt a single-step denoising Stable Diffusion, if the random noise is far from the target image in the latent space, this movement cannot be completed in a single step, making it impossible to fit the inverse prompt.
+Since we adopt a single-step denoising Stable Diffusion, if the input noise is far from the target image in the latent space, this movement cannot be completed in a single step, making it impossible to fit the inverse prompt.
 As shown in Figure 4(a), using random noise as input, after the loss value converges, the generated frame still has noticeable differences from the target frame, including blurring, noise artifacts, and inconsistent details.
-Therefore, we need to reduce the distance between the input noise and the target image in the latent space.
+Therefore, we need to reduce the distance between the input noise and the target frame in the latent space.
 We observe that in a video, adjacent frames are close in the latent space.
@@
 $\gamma$ is a hyperparameter that controls the degree of noise addition, which we set to $0.95$ in the experiment.
-$N^{t}$ is used as the noise input to Stable Diffusion for denoising the current frame.
+$N^{t}$ is used as the noise input to Stable Diffusion for generating the current frame.
 As shown in Figure 4(c), compared to random noise, the noisy previous frame can reduce the distance in the latent space, resulting in a generated frame that better matches the target frame.
@@
 Therefore, to reduce the overall error, reconstruction losses tend to abandon the fitting of edges and details, resulting in overly smooth and blurry images, as shown in Figure 4(b).
-To make the generated frames sharp and clear, one approach is to use perceptual loss instead of reconstruction loss, such as LPIPS [55].
+To make the generated frames sharp and clear, one approach is to use perceptual loss instead of reconstruction loss, such as LPIPS (zhang2018unreasonable, ).
 Perceptual loss is based on deep learning and can estimate the subjective quality of images as perceived by the human eye.
@@
 However, perceptual loss aims to maximize the overall subjective quality of the image without focusing on the exact consistency of each pixel, leading to misalignment between the generated and target images.
-Therefore, to simultaneously ensure pixel alignment and subjective quality, we combine the reconstruction and perceptual loss as the fitting loss $D$:
+To simultaneously ensure pixel alignment and subjective quality, we combine the reconstruction and perceptual loss as the fitting loss $D$:
 $D=\alpha*D_{rec}(x,x_{gt})+(1-\alpha)*D_{per}(x,x_{gt})$ (2)
@@
 $\alpha$ is a hyperparameter that balances pixel alignment and perceptual quality.
-In our experiments, we set $\alpha$ to 0.8.
+In our settings, we set $\alpha$ to 0.8.
 The final result is shown in Figure 4(c).
-It can be observed that by jointly optimizing the reconstruction loss and the perceptual loss, the images generated by Promptus ensure pixel-level alignment while maintaining a sharp appearance, making them almost identical to the ground truth images.
+It can be observed that by jointly optimizing the reconstruction loss and the perceptual loss, the image generated by Promptus ensures pixel-level alignment while maintaining a sharp appearance, making them almost identical to the ground truth images.
+Prompt Regularization.
+We found that even with real video as supervision, fitting occasionally fails.
+This is because as gradient descent progresses, the distribution of prompts gradually shifts away from the distribution of CLIP embeddings, resulting in an inability to generate meaningful images.
+To address this, we propose prompt regularization to constrain the distribution of prompts:
+$\lambda=\left\|\overline{c}-\mu\right\|_{2}$ (3)
+Here, $\mu$ is the mean of embeddings output by CLIP (which we obtained through statistics on a large amount of text, specifically $\mu=-0.168$). The final loss function $L$ is as follows:
+$L=\beta*D+(1-\beta)*\lambda$ (4)
+Where $\beta$ is a hyperparameter used to balance the fitting loss and the prompt regularization. In our experiments, we set it to 0.9.
 Figure 4: Visualization of prompt fitting results.
@@
 The results demonstrate that the noisy previous frame and the perceptual loss both contribute to the visual quality.
-3.2 Low-rank Decomposition based Prompt Bitrate Control
-According to §3.1, Promptus uses embeddings as prompts instead of text.
-However, embeddings are $m*n$ matrices (e.g., $1024*77$), where each element is a high-bit floating-point number (e.g., 32-bit float type), resulting in a much higher bitrate.
-To reduce the bitrate of prompts, Promptus has two directions: First, dimensionality reduction decreases the number of parameters in the prompt.
-Second, quantization reduces the number of bits for each parameter in the prompt.
+3.2 Low-rank Decomposition based Prompt Adaptive Bitrate Control
+According to §3.1, each prompt is an $m*n$ embedding matrix (e.g., $1024*77$), where each element is a floating-point number (e.g., 32-bit float), resulting in a fixed and high bitrate.
+To precisely control the bitrate of prompts, Promptus has two directions: First, dimensionality reduction decreases the number of prompt parameters.
+Second, quantization reduces the number of bits for each parameter.
 Low-rank matrix decomposition.
-To perform dimensionality reduction on the prompt, the most straightforward method is to first fit the complete prompt and then perform dimensionality reduction algorithms such as SVD (Singular Value Decomposition) or PCA (Principal Component Analysis) on it.
-However, these methods only perform dimensionality reduction based on the data distribution of the prompt, ignoring the impact of prompt degradation on the generation results of Stable Diffusion.
+To perform dimensionality reduction on the prompt, the most straightforward method is to first fit the complete prompt and then perform dimensionality reduction algorithms such as Singular Value Decomposition or Principal Component Analysis on it.
+However, these methods only perform dimensionality reduction based on the data distribution of the prompt, ignoring the impact of prompt degradation on the generation results.
 This inevitably leads to a degradation in the quality of the generated images.
 Therefore, instead of performing explicit dimensionality reduction, Promptus proposes to directly fit a low-dimensional prompt end-to-end, thereby reducing the quality degradation caused by dimensionality reduction.
-To achieve this, Promptus integrates the inverse process of CANDECOMP/PARAFAC decomposition [22] into gradient descent fitting.
+To achieve this, Promptus integrates the inverse process of CANDECOMP/PARAFAC decomposition (harshman1970foundations, ) into gradient descent fitting.
 Specifically, Promptus calculates the embedding $c$ as follows:
-$\vspace{-6mm}c=\frac{u*v}{\sqrt{r}}$ (3)
-where $u$ and $v$ are two low-rank factor matrices, and $r$ is the rank of the embedding.
+$c=\frac{u*v}{\sqrt{r}}$ (5)
+where $u$ and $v$ are two low-rank factor matrices with dimensions $m*r$ and $r*n$, respectively.
+$r$ is the rank of the embedding.
 $u$ and $v$ compose the embedding $c$ through outer product and normalization.
@@
 Compared to the embedding size of $m*n$ (e.g., $1024*77$), the total size of $u$ and $v$ is $(m+n)*r$.
-Therefore, reducing $r$ significantly lowers the bitrate.
+So reducing $r$ significantly lowers the bitrate.
 However, on the other hand, when Rank $r$ is smaller, the embedding is constrained to be a low-rank matrix, resulting in weaker representational capability and inability to fit high-frequency details in the image, as shown in Figure 7.
-Consequently, it is necessary to dynamically select the most appropriate $r$ based on the currently available network bandwidth to trade off between bitrate and quality.
+So it is necessary to dynamically select the most appropriate $r$ based on the current network bandwidth to trade off between bitrate and quality.
 Fitting-aware quantization.
 Although the number of parameters in the prompt has been significantly reduced through low-rank matrix decomposition, each parameter in $u$ and $v$ is still a high-bit floating-point number (such as 32-bit float type).
-Therefore, to further reduce the bitrate, it is necessary to quantize $u$ and $v$, reducing the number of bits for each parameter (such as lowering it to 8 bits).
-The most straightforward approach is to first fit $u$ and $v$ and then perform quantization.
-However, this post-quantization technique inevitably leads to quality loss.
-To address this, Promptus incorporates quantization into the fitting process, automatically compensating for the quantization loss through end-to-end gradient descent.
-We tested the impact of different quantization configurations on quality.
-Compared to traditional post-quantization, our method can reduce the number of bits from 32 to 8 with almost no quality loss.
-3.3 Prompt inter-frame compression based on temporal smoothing
+Therefore, to further reduce the bitrate, it is necessary to quantize $u$ and $v$, reducing the number of bits for each parameter.
+We adopt differentiable fake quantization (zafrir2019q8bert, ) and incorporate quantization into the fitting process, automatically compensating for the quantization loss through end-to-end gradient descent.
+We test the impact of different quantization configurations on quality.
+The results indicate our method reduces the number of bits from 32 to 8 without quality loss.
+3.3 Interpolation-aware Fitting based Prompt Inter-frame Compression
+Figure 5: The prompt interpolation not only fully preserves the details but also successfully fit the motion.
 Figure 6: At the receiver, Promptus approximates the unreceived frames by linear interpolation in the prompt space.
-According to §3.1, Promptus fits each frame of the video as an independent prompt, without considering the correlation between prompts across frames.
+Thus, instead of sending the prompt for each frame, only the prompts for keyframes need to be transmitted sparsely, thereby achieving inter-frame compression.
+According to §3.1, Promptus fits each frame of the video as an independent prompt, without considering the correlation across frames.
 Therefore, during video streaming, Promptus needs to transmit a prompt for each frame, resulting in a linear increase in bitrate as the frame rate rises.
@@
 The most straightforward solution is to reshape the prompt of each frame into a two-dimensional matrix and encode it using a video codec.
-However, unlike conventional videos, we found that this reshaped prompt looks like random noise, lacking any patterns, structures, or smooth regions, causing video codecs to no longer work.
-For inter-frame compression of prompts, our insight is: prompts are high-level semantics, so the prompts of continuous video frames should change continuously.
-If two temporally close frames are also sufficiently close in the prompt space, then the prompts of the frames between these two frames can be approximated by linear interpolation.
+However, unlike conventional videos, we found that this reshaped prompt looks like random noise, lacking any patterns, structures, or smooth regions, causing codecs to no longer work.
+For inter-frame compression of prompts, our insight is: prompts are high-level semantics, so the prompts of continuous video frames can change continuously.
 With this, during streaming, we only need to sparsely transmit the prompts of a few frames (as keyframes), and the prompts of the remaining frames can be obtained by linear interpolation of the keyframe prompts, as illustrated in Figure 6.
-Since only a small portion of keyframes need to transmit prompts, inter-frame compression is achieved.
-To ensure that adjacent frames are sufficiently close in the prompt space, we add temporal smoothing regularization to the embedding during fitting, as follows:
-$\lambda=\left\|c^{t}-c^{t-1}\right\|_{2}$ (4)
-Here, $c^{t}$ represents the embedding of the frame currently being fitted, while $c^{t-1}$ denotes the embedding of the previously fitted frame. The final loss function $L$ is as follows:
-$L=\beta*D+(1-\beta)*\lambda$ (5)
-Where $\beta$ is a hyperparameter used to balance the fitting loss and the temporal smoothing regularization. In our experiments, we set it to 0.2.
-Figure 6: Visualization of the prompt interpolation results.
-When applying temporal smoothing regularization, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video.
-Temporal smoothing regularization works, as shown in Figure 6.
-Without temporal smoothing regularization, the interpolation results suffer from severe distortions, disrupting the video’s content and structure.
-With temporal smoothing regularization, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video.
-4 Prompt Streaming System Design and Implementation
+Since keyframes account for only a small portion of the total frames, inter-frame compression is achieved.
+To ensure that the interpolated prompts can generate meaningful frames, we integrate prompt interpolation into the gradient descent fitting.
+Specifically, we randomly initialize prompts for keyframes, interpolate prompts for each intermediate frame, and generate frames.
+At this point, the prompts for intermediate frames exist as intermediate variables, so they are not optimized; instead, gradients are backpropagated to keyframes to directly optimize their prompts.
+Prompt interpolation works, as shown in Figure 5, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video.
+In practice, scene changes often occur within the video, at which point prompt interpolation no longer works.
+Therefore, we will continuously detect scene changes and treat the new scenes as new videos.
+For more details, please refer to §4.1.
+4 Prompt Streaming System Implementation
 §3 shows how Promptus inverts videos into pixel-aligned prompts.
-In this section, we will further describe how Promptus utilizes this Prompt Inversion technique to design and implement a streaming system.
+In this section, we will further describe how Promptus utilizes this Prompt Inversion technique to design a streaming system.
 Specifically, it will be divided into two parts: the sender side and the receiver side.
@@
 Note that this noise $Z^{0}$ will also be sent to the receiver side along with the prompt for generating the first frame image.
-On the other hand, due to the absence of a previous frame, Promptus does not apply temporal smoothing regularization for the Fitting of the first frame.
 Re-initialization for abrupt scene changes.
@@
 Sparse prompt streaming.
-As stated in §3.3, through temporal smoothing regularization, the prompts of most frames can be approximated by linear interpolation of the prompts of keyframes.
+As stated in §3.3, through interplation-aware fitting, the prompts of most frames can be approximated by linear interpolation of the prompts of keyframes.
 Therefore, Promptus defines the keyframe interval as $K$ (the choice of $K$ is discussed in §5.2), adding one keyframe every $K$ frames.
 Only the prompts of keyframes will be streamed to the receiver.
-Note that when the aforementioned abrupt scene changes occur, the difference between frames in the latent space is significant, and the interpolation of prompts no longer works.
-In this case, the last frame before the scene change will also become a keyframe and be sent.
+Note that when the aforementioned abrupt scene changes occur, the difference between frames in the prompt space is significant, and the interpolation of prompts no longer works.
+In this case, the last frame before the scene change will also become a keyframe.
 The frames after the scene change will start a new count.
@@
 Therefore, in prompt streaming, it is necessary to increase the prompt bitrate as much as possible while avoiding network congestion to maximize the user experience.
-Promptus adopts the WebRTC [1] framework for prompt streaming, which dynamically probes the currently available bandwidth.
-Consequently, Promptus adaptively adjusts the rank to make the prompt bitrate close to the probed bandwidth.
-It is worth noting that the prompts transmitted by Promptus are not encoded.
-Therefore, compared to codec-based streaming, Promptus can precisely control the bitrate.
+Consequently, Promptus fits multiple prompts of different ranks in advance.
+During streaming, Promptus selects the prompt with the bitrate closest to the estimated bandwidth for transmission.
+It is worth noting that the prompts transmitted are not entropy-encoded.
+Thus, compared to codec-based streaming, Promptus can precisely control the bitrate.
 4.2 The receiver side
@@
 Video generation based on sparse prompts.
-After receiving the prompt of keyframe $i$, the receiver first performs linear interpolation between the prompts of adjacent keyframes $i-k$ and $i$ to approximate the prompts of the intermediate $K-1$ frames.
+As shown in Figure 6, after receiving the prompt of keyframe $i$, the receiver first performs linear interpolation between the prompts of adjacent keyframes $i-k$ and $i$ to approximate the prompts of the intermediate $K-1$ frames.
 Afterward, these prompts are used for Stable Diffusion’s image generation.
-As described in §3.1, the generated $Z$ for each frame will be noised and used as the input noise for the generation of the next frame.
-Note that since the ratio of the noise (95%) in the noisy previous frame is much larger than the ratio of the previous frame (5%), the error or complete loss of the previous frame (earlier frames can be used) has almost no impact on the generation of subsequent frames.
+As described in §3.1, the generated $Z$ for each frame will be noised and used as the input noise for the next frame’s generation.
+Note that since the ratio of the noise (95%) in the noisy previous frame is much larger than the ratio of the previous frame (5%), the error of the previous frame has almost no impact on the generation of subsequent frames.
 Real-time video generation.
 The frame rate of video playback at the receiver depends on Promptus’s image generation speed, making real-time generation crucial.
-Promptus’s image generation consists of five parts: prompt dequantization, prompt composition, prompt interpolation, adding noise to previous frame and Stable Diffusion image generation.
+Specifically, Promptus’s generation consists of five parts: prompt dequantization, prompt composition, prompt interpolation, adding noise to previous frame and Stable Diffusion generation.
 The first four parts only involve simple linear calculations, so their time consumption can be ignored.
 The speed of Stable Diffusion becomes the bottleneck.
-To address this, we adopt the StreamDiffusion [28], which accelerates Stable Diffusion image generation to 100 FPS through batch processing.
-By this, Promptus achieves real-time video generation on the receiver.
+To address this, following StreamDiffusion (kodaira2023streamdiffusion, ), we first replace the VAE in Stale Diffusion with a more lightweight TAESD.
+To further accelerate generation, instead of using PyTorch, we adopt TensorRT to build inference engines for Promptus.
+In this way, Promptus achieves real-time video generation at the receiver.
+A detailed evaluation of generation speed can be found in §5.9.
 5 Evaluation
-Our four main evaluation results are as follows: First, Promptus achieves scalable bitrate, and its quality advantage over the baselines becomes more significant as the target bitrate decreases.
-Second, Promptus is general to different video domains, and the more complex the video content, the greater the quality advantage of Promptus compared to the baselines.
-Third, under real-world network traces, Promptus enhances the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%.
-Fourth, Promptus can generate videos from prompts in real-time at a speed exceeding 150 FPS.
 5.1 Experiment Setup
-Test videos: To validate the generalizability of Promptus across different video domains, we selected 7 videos from 4 datasets with vastly different content, as summarized in Table 1.
-Specifically, the domains of these videos span natural landscapes and human activities, outdoor long-range scenes and indoor close-up scenes, real-world scenes and CG-synthesized scenes, 3D gaming scenes and 2D animations.
-All videos are cropped and resized to a resolution of 512*512, with a frame rate of 30 FPS.
-Since Promptus uses a pre-trained and frozen Stable Diffusion, it does not involve model training and does not require training videos.
-However, for fairness, we still checked that the above test videos were not used during the pre-training of Stable Diffusion.
-Table 1: Test videos summary
-Dataset #Videos #frames Description
-QST [53] 2 300 Natural landscapes, outdoor distant view
-UVG [32] 2 300 Human activity, face, hand, indoor close-up
-GTA-IM [17] 2 300 3D Game recording, CG-synthesized scenes
-Animerun [43] 1 60 2D animation, cartoon
-Total 7 960
-Baselines: We compare Promptus with two baselines: H.265 [5] and VAE [27, 38].
-H.265 is an advanced traditional codec that achieves compression by utilizing hand-designed intra-frame prediction, motion compensation, and transform coding techniques.
-Moreover, H.265 can control the bitrate of the encoded video by adjusting the quantization parameters, balancing quality and bitrate.
-Since it does not involve learnable parameters, it has good generality.
-VAE is a deep learning-based neural codec.
+Test videos: To validate the generality across different domains, we test Promptus on four video datasets with vastly different content: QST zhang2020dtvnet , UVG mercat2020uvg , GTA-IM caoHMP2020 and Animerun siyao2022animerun .
+The domains of these videos span natural landscapes and human activities, outdoor long-range scenes and indoor close-up scenes, real-world scenes and CG-synthesized scenes, 3D video games and 2D animations.
+For consistency, videos are cropped and resized to a resolution of 512*512, with a frame rate of 30 FPS.
+Note that Promptus supports flexible resolutions and we test other resolutions in §5.6.
+Baselines: We compare Promptus with four baselines: H.265 (265, ), VAE (rombach2022high, ; kingma2014auto, ), NAS (yeo2018neural, ) and VQGAN (esser2021taming, ; li2023reparo, ):
+  • H.265 is an advanced traditional codecs that achieve compression by utilizing hand-designed modules.
+  • VAE is a learning-based neural codec.
 Both its encoder and decoder are trainable neural networks.
-The encoder maps the input image to a low-dimensional latent variable, while the decoder reconstructs the image from the latent variable.
-The dimensionality of the latent variable is usually much smaller than that of the original image, thus achieving compression.
-For inter-frame compression of VAE, we first quantize the latent variables of each frame and then encode them into a video using H.265.
-The quantization process causes almost no quality degradation, and the quality degradation is mainly due to video encoding.
-Therefore, we adjust the target bitrate of video encoding to balance the quality and bitrate of VAE.
-For fairness, the training set of VAE is the same as that of Stable Diffusion, both consisting of 5.85 billion images [6].
-Metric: To evaluate video quality, we adopt the LPIPS [55] (Learned Perceptual Image Patch Similarity) instead of the traditional SSIM and PSNR.
-This is because LPIPS has a significantly higher correlation with human subjective ratings compared to SSIM and PSNR [51, 55], which means that LPIPS better reflects human subjective perception of video quality.
-LPIPS uses a pre-trained neural network to extract layer-wise features from the original image and the distorted image, and then calculates the normalized L2 distance between the features as the LPIPS value.
-Therefore, a smaller LPIPS value indicates a higher quality of the distorted image.
-Specifically, we use VGG [41] pre-trained on ImageNet [18] as the feature extractor for LPIPS.
-Network trace: We collected 6 network traces from real-world scenarios such as subways, driving, and walking, under 2G, 3G, and 4G networks.
-The traces consisted of 1 from a 4G network, 2 from 3G, and 3 from 2G.
-Each trace lasted between 5 and 30 seconds, totaling over 100 seconds.
-Our traces aimed to test various weak network conditions such as poor signal coverage, network overload, high-speed movement, and frequent switching of mobile communication networks.
-The average bandwidth per second ranged between 50 kbps and 4000 kbps, while the one-way network delay was approximately 30 ms to 100 ms.
-Testbed: To test the streaming performance, we use WebRTC [1] to deploy Promptus and the baselines.
-WebRTC is a real-time video communication framework that dynamically probes the currently available bandwidth to avoid congestion.
-To simulate real network conditions, we use Mahimahi [34] to replay the above network traces.
+The encoder maps the input image to low-dimensional latent variables, while the decoder reconstructs the image from the latent variables.
+The dimensionality of the latent variables is usually much smaller than that of the original image, thus achieving compression.
+For inter-frame compression, we encode the latent variables using H.265.
+  • NAS represents neural-enhanced streaming, which trains overfitted super-resolution DNNs for each video.
+These DNNs are transmitted along with low-resolution videos.
+At the receiving end, DNNs are employed to enhance the quality of low-resolution videos.
+To control the total bitrate, we adjust the bitrate of the low-resolution video and the number of DNN parameters transmitted.
+  • VQGAN stands for generative streaming.
+To compress the video, it first divides the video into patches, maps each patch to a latent variable using a VAE, and then quantizes each latent variable into a 10-bit token index using a codebook.
+To support general videos, the codebook is trained on diverse visual domains plummer2015flickr30k , not limited to facial videos (like Reparo (li2023reparo, )).
+At the receiver, the codebook is shared, so the token indices can be reconstructed into the video.
+To control the bitrate, we adjust the number of patches (token indices) for streaming.
+Figure 7: Visualization of the fitting results at different ranks.
+It can be seen that as the rank increases, the prompt can fit more details.
+For example, when the rank is 4, the earrings in (a) are lost, and the lamp’s color in (d) is inconsistent with the ground truth.
+When the rank increases to 16, the earrings in (b) are successfully fitted, and the lamp’s color in (e) is also corrected.
+Metric: To evaluate video quality, we mainly adopt the LPIPS (zhang2018unreasonable, ), because Promptus is more oriented towards the semantic communication task rather than video compression.
+Our primary goal is to keep the semantic information correct at low bitrates.
+LPIPS better reflects human subjective perception quality (yu2023language, ; zhang2018unreasonable, ), making it more suitable for evaluating semantic distortion.
+Smaller LPIPS indicates higher quality.
+We also evaluate traditional PSNR and SSIM as objective metrics, with discussions provided in §6.
+Network trace: We collect network traces from real-world scenarios such as subways, driving, and walking, under 2G, 3G, and 4G networks.
+Our traces aim to test various weak network conditions such as poor signal coverage, network overload, high-speed movement, and frequent switching of mobile wireless networks.
+The average bandwidth per second ranges between 50 kbps and 4000 kbps, while the one-way network delay is about 30 ms to 100 ms.
+We use Mahimahi (netravali2015mahimahi, ) to replay the these network traces.
 The queue length of Mahimahi is set to 60, and the drop-tail strategy is adopted.
@@
 This section demonstrates how different parameter configurations affect the quality-bitrate tradeoff of Promptus.
-Figure 7: Visualization of the fitting results at different ranks.
-It can be seen that as the rank increases, the prompt can fit more details.
-When the rank is 4, the earrings are lost.
-When the rank increases to 16, the earrings are successfully fitted.
 Prompt rank.
@@
 This is because the larger the rank, the stronger the representational capability of the prompt, allowing it to fit the target image more accurately, as described in §3.2.
-We present a visualization of the fitting results at different ranks, as shown in Figure 7.
-It can be seen that as the rank increases, the prompt can fit more details.
-For instance, when the rank is 4, the earrings are lost in the fitting result, while when the rank increases to 16, the earrings are successfully fitted.
+We present a visualization of the fitting results at different ranks in Figure 7.
+It indicates that as the rank increases, the prompt can fit more details.
+For instance, when the rank is 4, the earrings are lost.
+When the rank increases to 16, the earrings are successfully fitted.
 Keyframe interval.
@@
 For example, when the prompt rank is 16, reducing the keyframe interval from 10 to 1 decreases the LPIPS from 0.274 to 0.221.
-This is because, when the interval is smaller, the distance between keyframes in the prompt space is smaller.
-Since we constrain the prompt space to be temporally smooth, the linear interpolation of keyframe prompts can more accurately approximate the prompts of intermediate frames, as described in §3.3.
-Figure 8: The impact of prompt rank and keyframe interval on visual quality. It indicates that visual quality improves with increasing prompt rank and decreasing keyframe interval.
+This is because, when the keyframe interval is smaller, the distance between keyframes in the prompt space is smaller.
+At this time, the linear interpolation of keyframe prompts can more accurately approximate the prompts of intermediate frames, as described in §3.3.
+Figure 8: The impact of prompt rank and keyframe interval on visual quality. It shows that visual quality improves with increasing rank and decreasing keyframe interval.
+Figure 9: Tradeoff between bitrate and quality for Promptus. The quality monotonically increases with the increase in bitrate. Besides, Promptus can achieve scalable bitrates.
+Figure 10: Average frame quality on 4 datasets. It proves the generality of Promptus across domains. The more high-fidelity details a video has, the greater the advantage of Promptus.
+(a) 140 kbps
+(b) 280 kbps
+(c) 360 kbps
+(d) 540 kbps
+Figure 11: Frame quality CDF and mean of Promptus and baselines at four target bitrates.
+It indicates Promptus achieves better transmission efficiency across all bitrate levels.
+The lower the bitrate, the greater the advantage of Promptus.
+Detail in  §5.3.
 Quality-bitrate tradeoff.
 Increasing the prompt rank and reducing the keyframe interval lead to a rapid rise in bitrate while improving quality.
-Therefore, we present the tradeoff between bitrate and quality for Promptus, as shown in Figure 9.
+Therefore, we present the tradeoff between bitrate and quality, as shown in Figure 10.
 First, it illustrates that overall, the quality monotonically increases with the increase in bitrate.
@@
 Third, at the same bitrate, the quality varies for different configurations.
-For instance, when the bitrate is 550 kbps, the configuration with a rank of 8 and a keyframe interval of 4 has an LPIPS of 0.255, while the configuration with a rank of 16 and a keyframe interval of 8 has an LPIPS of 0.264, which is lower in quality than the former.
-So at the same bitrate, Promptus tends to choose configurations with smaller keyframe intervals for streaming.
-Figure 9: The tradeoff between bitrate and quality for Promptus. The quality monotonically increases with the increase in bitrate. Besides, Promptus can achieve scalable bitrates.
-5.3 Compression efficiency
-This section demonstrates the compression efficiency of Promptus.
-Figure 10 shows the CDF of the frame quality for Promptus and baselines under 4 bitrate levels.
-Since a lower LPIPS represents better visual quality, a leftward shift of the curve in the figure represents more high-quality frames, and thus higher compression efficiency.
+For instance, when the bitrate is 550 kbps, the parameter configuration with a rank of 8 and a keyframe interval of 4 has an LPIPS of 0.255, while the parameter configuration with a rank of 16 and a keyframe interval of 8 has an LPIPS of 0.264, which is lower in quality than the former.
+So at the same bitrate, we tend to choose parameter configurations with smaller keyframe intervals.
+5.3 Transmission efficiency
+This section demonstrates the transmission efficiency.
+Figure 11 shows the CDF of the frame quality under 4 bitrate levels.
+Since a lower LPIPS represents better visual quality, a leftward shift of the curve in the figure represents more high-quality frames, and thus higher transmission efficiency.
 We also calculate the average LPIPS for each method, represented by the vertical lines in the figure.
-(a) 140 kbps
-(b) 280 kbps
-(c) 360 kbps
-(d) 540 kbps
-Figure 10: Frame quality CDF and mean of Promptus and baselines at four target bitrates.
-It indicates Promptus achieves better compression efficiency across all bitrate levels.
-Besides, the lower the bitrate, the greater the advantage of Promptus.
-First, Promptus achieves better compression efficiency across all bitrate levels.
+First, Promptus achieves better transmission efficiency across all bitrate levels.
 For example, in Figure 11, the curves of Promptus are all to the left of the baseline curves.
-Second, the lower the bitrate, the greater the advantage of Promptus.
+Second, Promptus can achieve more than a 4x bandwidth reduction while preserving the same perceptual quality.
+For example, the average LPIPS of Promptus under 140 kbps is better than H.265 under 540 kbps.
+Third, the lower the bitrate, the greater the advantage of Promptus.
 At a high level, as the bitrate decreases from 540 kbps to 140 kbps, the distance between Promptus’s curves and the baselines’ curves gradually widens.
-At a low level, when the bitrate is 540 kbps, the average LPIPS of Promptus is 0.018 and 0.085 lower than VAE and H.265, respectively.
-When the bitrate is 140 kbps, this difference further increases to 0.139 and 0.118.
-This is because when the bitrate is reduced, H.265 uses coarser quantization and loses many high-frequency details, resulting in blurriness and block artifacts in the video, which significantly impairs the perceptual quality.
-VAE mitigates this phenomenon by mapping images to a low-dimensional latent space, reserving more bitrate for video coding.
+At a low level, when the bitrate is 540 kbps, the average LPIPS of Promptus is 0.018, 0.085 and 0.037 lower than VAE, H.265 and VQGAN, respectively.
+When the bitrate is 140 kbps, this difference further increases to 0.139, 0.118 and 0.046.
+This is because when the bitrate is reduced, the traditional codecs use coarser quantization and lose many high-frequency details, resulting in blurriness and block artifacts in the video, which significantly impairs the perceptual quality.
+VAE mitigates this phenomenon by mapping images to a low-dimensional latent space, reserving more bitrate for video encoding.
 Therefore, at most bitrates, the quality of VAE is superior to H.265.
@@
 This prevents Promptus from accurately describing the video content, resulting in slight misalignments in the generated frames.
-However, thanks to the inherent image generation capability of Stable Diffusion, Promptus’s frames still have good sharpness and details, and thus the perceptual quality is better than VAE and H.265.
-5.4 Generality
-This section demonstrates the generality of Promptus across different domains.
-Figure 12 shows the mean of the frame quality for Promptus and two baselines on four different datasets (as described in Table 1).
+However, thanks to the inherent image generation capability of Stable Diffusion, Promptus’s frames still have good sharpness and details, thus outperforming the baselines.
+When the bitrate decreases, VQGAN uses fewer tokens, which also reduces its representational capability.
+However, it does not have gradient descent for quality compensation, making it inferior to Promptus.
+5.4 Generality on different domains
+This section proves the generality of Promptus across different domains. Figure 10 shows the average frame quality for Promptus and four baselines on four datasets.
+First, Promptus has good transmission efficiency on different domains.
+As shown in Figure 10, Promptus achieves lower average LPIPS compared to most baselines on each dataset.
+To intuitively demonstrate this gain, we visualize the compression results of each method on the four datasets at low bitrates (such as 225 kbps) in Figure 12.
+It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
 Figure 12: Visualization of the compression results on different datasets.
-It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
-Figure 12: Mean frame quality of Promptus and baselines on 4 different datasets.
-It demonstrates the generality of Promptus across different domains.
-Besides, the more high-frequency details a video has, the greater the advantage of Promptus.
-First, Promptus achieves better compression efficiency on each dataset.
-This is because, as shown in Figure 12, Promptus achieves lower average LPIPS compared to the baselines on each dataset.
-To intuitively demonstrate this improvement, we also visualize the compression results of the three methods on the four datasets at lower bitrates (such as 225 kbps), as shown in Figure 11.
-It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
+It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details.
 Second, the more high-frequency details a video has, the greater the advantage of Promptus.
@@
 This is because for 2D animations with large areas of solid colors and simple details, H.265’s intra-frame prediction, block partitioning, and motion compensation techniques can handle them well, so Promptus’s performance gain is small.
-For detail-rich real-world videos, H.265 discards more high-frequency information during compression, thus damaging the perceptual quality (especially at lower bitrates).
-Although Promptus also loses high-frequency information from the original image, Stable Diffusion completes the lost high-frequency information based on prior knowledge during generation, resulting in a smaller perceptual quality loss.
+This is also why Promptus performs worse than NAS on this dataset, as simple 2D animations are the domain where super-resolution performs well.
+For detail-rich real-world videos, H.265 discards more high-frequency information during compression, thus damaging the perceptual quality.
+Although Promptus also loses high-frequency information, Stable Diffusion completes some lost information based on prior knowledge during generation, resulting in a smaller quality loss.
+Figure 13: Fitting results for challenging examples, with a bitrate of 225 kbps.
+Although the Stable Diffusion itself struggles to generate specific elements, such as fingers and text, Promptus is able to fit them well.
+This is because end-to-end gradient descent fitting can compensate for the limitations of the Stable Diffusion itself.
+Since Stable Diffusion itself struggles to generate specific elements, such as fingers and text.
+To further demonstrate the generality of Promptus, we also show its fitting results for these challenging examples in Figure 13.
+The results indicate that Promptus is able to fit them well.
+This is because end-to-end gradient descent fitting can compensate for the limitations of the Stable Diffusion itself.
+5.5 Temporal Consistency
+This section evaluates the temporal consistency of the generated videos.
+As shown in Figure 14, we visualize the entire videos as space-time X-t slices (li2024generative, ), with three scanning lines at different heights marked on the left.
+Both of our method and H.265 are at a bitrate of 225 kbps.
+Results demonstrate that although Prompus generates video frame by frame, it maintains good temporal consistency.
+First, it can be seen that our videos are temporally aligned with the ground truth videos in terms of motion.
+Second, our videos exhibit almost no flickering phenomena.
+For example, pixels in the X-t slices mostly move smoothly along the t dimension.
+This is because the proposed interpolation-aware fitting (§3.3) allows each frame to be as similar as possible to the ground truth, and video deflickering lei2023blind is also used to further alleviate inconsistency between frames.
+Third, compared to H.265, our method reduces visual degradation caused by blocking artifacts.
+As shown in the red box of Figure 14, H.265 exhibits some zig-zag phenomena in X-t slices, while our method is free from these artifacts.
+We provide some demo videos in our open-source repository to intuitively showcase the quality.
+Figure 14: To evaluate the temporal consistency of the generated videos, we show the videos second by second.
+We also visualize the entire videos as space-time X-t slices (li2024generative, ), with three scanning lines at different heights marked on the left.
+The results indicate that our videos are temporally aligned with the ground truth videos in terms of motion, and exhibit almost no flickering phenomena.
+More details are in §5.5.
+5.6 Flexible Resolution
+Figure 15: Promptus supports flexible resolutions (§5.6).
+This section shows that Promptus supports flexible resolutions.
+As illustrated in Figure 15, our generated images exhibit near-perfect visual alignment with ground truth across all resolutions.
+In fact, the resolutions supported by Promptus depend on the generator employed.
+Since Stable Diffusion supports flexible resolutions, Promptus naturally inherits this characteristic.
+This also explains a potential limitation: Promptus performs poorly on low-resolution videos because Stable Diffusion cannot generate low-resolution images sdturbo .
+For instance, as shown in Figure 15, videos with 256*256 resolution achieved the lowest PSNR.
 5.7 Performance on real-world traces
-This section demonstrates the performance of Promptus under real network traces.
-We use the traces and testbed described in §5.1 to run Promptus and two baselines.
-Figure 13 shows the CDF and mean of the frame quality.
-Since the total length of the traces is greater than the total length of the test videos, we loop the test videos to run through the entire traces.
 Figure 16: The performance of Promptus under real network traces.
 On one hand, Promptus’s quality is overall higher than the baselines.
-On the other hand, Promptus can significantly reduce the ratio of severely distorted frames.
+On the other hand, Promptus can significantly reduce the ratio of severely distorted frames (whose LPIPS is higher than 0.32).
+This section demonstrates the performance of Promptus under real-world network traces.
+Since the total length of the network traces does not match the total length of the test videos, we loop the videos to align with the traces.
+Figure 16 shows the CDF and mean of the frame quality.
 First, Promptus’s quality is overall higher than the baselines.
-For example, the mean LPIPS of Promptus is 0.111 and 0.092 lower than VAE and H.265, respectively.
+For example, the mean LPIPS of Promptus is 0.111, 0.092, 0.047 and 0.063 lower than VAE, H.265, VQGAN and NAS, respectively.
 Second, Promptus can significantly reduce the ratio of severely distorted frames.
-For instance, only 5.2% of frames in Promptus have LPIPS higher than 0.32, while VAE and H.265 have 94.5% and 96.9%, respectively.
-These improvements are partly due to Promptus’s excellent compression efficiency, enabling it to provide higher perceptual quality at the same bitrate.
-On the other hand, since Promptus sends raw prompts without encoding (such as entropy coding), it can precisely control the target, thus making full use of bandwidth.
+For instance, only 5.2% of frames in Promptus have LPIPS higher than 0.32, while VAE, H.265, VQGAN and NAS have 94.5%, 96.9%, 53.2% and 98.2% respectively.
+These improvements are partly due to Promptus’s excellent transmission efficiency, enabling it to provide higher perceptual quality at the same bitrate.
+On the other hand, since Promptus sends raw prompts without entropy coding, it can precisely control the target bitrate, thus making full use of bandwidth.
+5.8 Ablation study of interpolation methods and generators
+Figure 17: Comparison of different interpolation methods.
+Each method is based on frame 0 and frame 10 to interpolate frames 1 to 9.
+The results demonstrate that latent interpolation fails to preserve the motion between frames, resulting in spatial overlaps and ghosting.
+This is because the adjacent frames in latent space are not linearly continuous, making the interpolation unreasonable.
+While pixel interpolation (huang2022rife, ) can preserve motion, it inevitably leads to noticeable artifacts in cases of occlusion, edges, or newly appearing objects due to incorrect matching, as illustrated by the green and red boxes in the figure.
+Our prompt interpolation successfully preserves motion while avoiding artifacts and keeping the edges sharp.
+Details can be found in §5.8
+In this section, we compare different interpolation methods and analyze why it should be prompt interpolation rather than latent or pixel interpolation.
+We also demonstrate why it is necessary to use the complete Stable Diffusion as the generator for inversion rather than just the VAE (Decoder).
+Specifically, for latent interpolation, we remove the U-Net from the inversion framework (described in §3), retaining only the VAE Decoder.
+Thus, we randomly initialize the frames in the latent space and apply gradient descent for fitting.
+After finishing the fitting, we perform linear interpolation on the fitting results (latent variables) of the key frames to approximate the intermediate frames, just like in prompt interpolation.
+The results are shown in "End-to-end latent interpolation" in Figure 17.
+For pixel interpolation, we apply RIFE (huang2022rife, ) (a real-time video frame interpolation algorithm) to the ground truth images of the key frames, directly estimating the intermediate frames.
+The results are shown in "End-to-end pixel interpolation" in Figure 17.
+Additionally, to compare with Promptus based solely on interpolation methods, we also conduct experiments where key frames (both the latent variables and images) come from Promptus, represented as "Latent interpolation" and "Pixel interpolation" shown in Figure 17.
+Each of the above methods is based on frame 0 and frame 10 to interpolate frames 1 to 9.
+First, the results demonstrate that latent interpolation fails to preserve the motion between frames, resulting in spatial overlaps and ghosting, as shown in the second and third rows of the Figure 17.
+This is because the adjacent frames in latent space are not linearly continuous, making the interpolation between frames unreasonable.
+This indicates that the VAE alone cannot serve as the generator for video inversion, since inter-frame compression does not work, and frame-by-frame transmission would result in significant bandwidth wastage.
+A feasible solution is to use codecs to encode the latent variables of each frame.
+However, as shown in Figure 11 and Figure 10, this solution performs worse than prompt interpolation.
+Second, while pixel interpolation can preserve motion, it inevitably leads to noticeable artifacts in cases of occlusion, edges, or newly appearing objects due to incorrect matching, as illustrated by the green and red boxes in the Figure 17.
+Third, our prompt interpolation successfully preserves motion while avoiding artifacts and keeping the edges sharp.
+This is because the U-Net converts frames from latent space to prompt space, where the frames are linearly continuous, allowing the interpolation to approximate motion.
+It proves that the U-Net (diffusion process) plays a significant role in compression performance.
 5.9 Overhead
-In this section, we analyze the overhead of Promptus’s generation.
+Figure 18: Visualization of fitting results at different iterations.
+The results indicate that for the first frame of a new video (or new scene), the fitting process takes thousands of iterations to converge since it starts from scratch.
+For subsequent frames, thanks to the continuous prompt space (§3.3) providing sufficient scene priors, the fitting process can converge in only a few hundred iterations.
+For example, the second frame in the figure requires only 120 iterations to achieve a visual quality comparable to that of 2000 iterations for the first frame.
+In this section, we analyze the overhead of Promptus.
 We conduct tests on an Nvidia 4090D GPU, using CUDA to accelerate Promptus.
 The resolution of the generated video is 512*512, and the rank of the prompt is 8.
-Table 2: Overhead of generating a frame
-Steps Time (ms) Memory (MB)
-Prompt dequantization 0.016 -
-Prompt composition 0.025 -
-Prompt interpolation 0.013 -
-Noised previous frame 0.012 -
-Stable Diffusion generation 6.160 -
-Total 6.226 8952
+The Stable Diffusion model has a total of 867M parameters.
+With more parameters (such as SDXL Turbo (sdxlturbo, )), the Stable Diffusion model has stronger generative ability, making prompt fitting easier and allowing for a higher compression ratio.
+However, this also leads to increased overhead in terms of memory usage and run time.
+Thus, the current version of Promptus employs SD 2.1 Turbo, which is relatively lightweight.
+Table 1: Time overhead of generating a frame §5.9.
+Steps Time (ms)
+Dequantization 0.016
+Prompt composition 0.025
+Prompt interpolation 0.013
+Noised frame 0.012
+Stable Diffusion generation 6.160
+Total 6.226
+Generation overhead.
 Table 1 shows the fine-grained overhead of each step in Promptus’s image generation.
-Specifically, from receiving a prompt to generating a frame, Promptus includes the following steps: prompt dequantization, prompt composition, prompt interpolation, adding noise to the previous frame, and Stable Diffusion image generation.
+Specifically, from receiving a prompt to generating a frame, Promptus includes the following steps: prompt dequantization, prompt composition, prompt interpolation, adding noise to the previous frame, and Stable Diffusion generation.
 Among them, most steps only involve simple linear computations, so the time overhead is almost negligible.
-In contrast, Stable Diffusion image generation accounts for the vast majority of the time overhead.
-Therefore, we adopt StreamDiffusion [28], which accelerates Stable Diffusion to real-time through batch processing techniques and implementation optimizations.
-The difference is that compared to the original StreamDiffusion, Promptus removes the CLIP module that converts text to embeddings, thus further accelerating image generation to 6.160 ms.
-In summary, the total time overhead of Promptus’s image generation is 6.226 ms, achieving real-time.
+In contrast, Stable Diffusion generation accounts for the vast majority of the time overhead.
+Therefore, following StreamDiffusion (kodaira2023streamdiffusion, ), we use TAESD and TensorRT to accelerate the Stable Diffusion to 6.16 ms.
+In summary, the total time overhead of Promptus’s image generation at the receiver is 6.226 ms, achieving real-time.
+Besides, the total memory usage is 7846 MB.
+Inversion overhead.
+The time overhead of inversion depends on two factors: the number of iterations needed for convergence and the time taken for each iteration.
+Among them, the time taken for each iteration is 150 ms.
+For the number of iterations, there is a significant difference between different frames.
+For the first frame of a new video (or new scene), the inversion takes thousands of iterations to converge since it starts from scratch.
+For subsequent frames, thanks to the continuous prompt space §3.3 providing sufficient scene priors, the inversion can converge in only a few hundred iterations.
+We present an example in Figure 18.
+It can be seen that the second frame requires only 120 iterations to achieve a visual quality comparable to that of 2000 iterations for the first frame.
+In practice, we set the number of iterations for the first frame to 10,000, and for subsequent frames to 500.
+As for memory usage, inversion occupies a total of 10 GB.
 6 Limitation
@@
 This is because prompt fitting requires iterative gradient descent.
-Although a single iteration is fast (on the same order of magnitude as the time overhead of real-time generation), the total time overhead is high due to the large number of iterations required for convergence (such as 500 iterations).
+Although a single iteration is fast, the total time overhead is high due to the large number of iterations required for convergence (such as 500 iterations).
+As a result, Promptus can currently only be used for Video on Demand and is not applicable for Live Video or Real-Time Communication.
 To address this, using more efficient gradient descent algorithms to reduce the number of iterations required for convergence is expected to accelerate prompt fitting to real-time.
@@
 This means that if an intermediate frame needs to be generated and played, it is necessary to wait until the subsequent keyframe is received, which introduces additional latency.
+Although this latency has little impact on Video on Demand, as DASH (sodagar2011mpeg, ) provides ample buffering to cover the interpolation latency.
+However, it is not suitable for latency-sensitive applications, such as WebRTC-based (razor, ) video conferencing and cloud gaming wu2023zgaming , where buffer sizes are small.
 To address this issue, designing keyframe extrapolation algorithms to replace interpolation is a future research direction.
@@
 A solution is to reduce the keyframe interval and send keyframes more densely.
-However, there also exist smoothly changing segments in the video, where densely sending keyframes brings little improvement to quality, thus wasting bandwidth.
+But there also exist smoothly changing segments in the video, where densely sending keyframes brings little improvement to quality, thus wasting bandwidth.
 In summary, sending keyframes uniformly in this paper is inefficient.
-In the future, adaptive keyframe needs to be designed to transmit densely in rapidly changing segments and sparsely in smoothly changing segments.
+In the future, adaptive keyframe needs to be designed to transmit densely in rapidly changing segments and sparsely in smooth segments.
+Performance on objective metrics.
+Although Promptus achieves significant improvements in perceptual quality, it does not perform equally well on some objective metrics (such as PSNR, SSIM), as shown in Table 2.
+This is because Promptus is more sensitive to sharp textures like all AIGC models, while PSNR and SSIM are more sensitive to smooth textures with large areas.
+As a result, the images generated by Promptus have sharper details with better perceptual quality, while the baselines are more blurred but with higher PSNR.
+Sometimes, the more details presented by Promptus, the lower the PSNR, while the more blurred the baselines are, the higher the PSNR can be.
+This phenomenon is also illustrated in the subjective examples in Figure 12.
+Table 2: Comparison of PSNR, SSIM, and LPIPS metrics on the UVG (mercat2020uvg, ) at a bitrate of 225 kbps.
+H.265 VAE NAS VQGAN Ours
+PSNR $\uparrow$ 32.685 28.008 33.466 25.740 30.378
+SSIM $\uparrow$ 0.892 0.646 0.780 0.648 0.702
+LPIPS $\downarrow$ 0.386 0.382 0.357 0.335 0.244
+Although PSNR and SSIM are widely used in the video compression task.
+However, in this paper, Promptus is more oriented towards the semantic communication task.
+The main goal is to keep the semantic information correct at low bitrates.
+LPIPS considers perceptual quality, making it more suitable for evaluating semantic distortion.
+After all, Promptus is a novel paradigm for high-fidelity semantic communications.
 7 Conclusion
-In this paper, we propose Promptus, a novel system that replaces video streaming with prompt streaming by inverting video frames into prompts for Stable Diffusion.
+In this paper, we propose Promptus, a novel system that replaces video streaming with prompt streaming by representing video frames as Stable Diffusion prompts.
 To ensure pixel alignment, a gradient descent-based prompt fitting framework is proposed.
-To achieve adaptive bitrate for prompts, a low-rank decomposition-based bitrate control algorithm is introduced.
-For inter-frame compression of prompts, a temporal smoothing-based prompt interpolation algorithm is proposed.
-Evaluations across various video domains and real network traces demonstrate Promptus can enhance the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%.
-Moreover, Promptus achieves real-time video generation from prompts at over 150 FPS.
-By pioneering the replacement of video codecs with prompt inversion and introducing prompt streaming, Promptus revolutionizes efficient video communication, surpassing the limitations of traditional approaches and paving the way for a new era of video streaming.
+To achieve adaptive bitrate, a low-rank decomposition-based bitrate control algorithm is introduced.
+For inter-frame compression, an interpolation-aware fitting algorithm is proposed.
+Evaluations across various video domains and real network traces demonstrate that, compared to H.265, Promptus can achieve more than a 4x bandwidth reduction while preserving the same perceptual quality.
+Promptus extends the boundaries of AIGC to video streaming, offering a semantic video communication paradigm.
+Promptus is now open-source, including real-time demos, generation engines, pre-fitted prompts, and comprehensive documentation: https://github.com/JiangkaiWu/Promptus.